### v1

# Relational Self-supervised Distillation with Compact Descriptors for Image Copy Detection

Abstract.
This paper addresses image copy detection, a task in online sharing platforms for copyright protection. While previous approaches have performed exceptionally well, the large size of their networks and descriptors remains a significant disadvantage, complicating their practical application. In this paper, we propose a novel method that achieves a competitive performance by using a lightweight network and compact descriptors. By utilizing relational self-supervised distillation to transfer knowledge from a large network to a small network, we enable the training of lightweight networks with a small descriptor size. Our approach, which we call Relational self-supervised Distillation with Compact Descriptors (RDCD), introduces relational self-supervised distillation (RSD) for flexible representation in a smaller feature space and applies contrastive learning with a hard negative (HN) loss to prevent dimensional collapse. We demonstrate the effectiveness of our method using the DISC2021, Copydays, and NDEC benchmark datasets, with which our lightweight network with compact descriptors achieves a competitive performance. For the DISC2021 benchmark, ResNet-50/EfficientNet-B0 are used as a teacher and student respectively, the micro average precision improved by 5.0%/4.9%/5.9% for 64/128/256 descriptor sizes compared to the baseline method.
Knowledge Distillation, Self-supervised Learning, Lightweight Network, Image Copy Detection
1. Introduction
Figure 1. Comparison of RDCD(Ours) and other image copy detection methods. RDCD utilizes a lightweight network and achieves high performance despite its compact descriptor sizes.
Visual copy detection is a task that determines whether two pieces of multimedia items originate from a single source. This is widely used in online sharing platforms, including social networks, to filter content for copyright protection. Image copy detection (ICD), a part of visual copy detection, is the task of detecting copied images from an existing database, where each image instance is treated as an individual category such as instance matching problem (Douze et al., 2021a). This task has also been explored under the name of near-duplicate image detection (Chum et al., 2008) or content-based image retrieval (Kim, 2003). In this context, their approaches are centered on category-level recognition or instance-level recognition, focusing on identifying broad categories or specific objects, respectively. However, ICD addresses a more nuanced challenge by focusing on exact image copies that often involve severe transformations such as re-encoding, resizing, merging, cropping, warping, or color distortion. Traditional approaches in this domain have primarily employed either global descriptors (Kim, 2003; Wang and Jiang, 2015) or local descriptors such as bag-of-words (Zhou et al., 2010), which may struggle with such transformations. To better address these challenges, recent advancements have leveraged self-supervised learning (SSL) methods such as DINO (Caron et al., 2021), SimCLR (Chen et al., 2020b), MoCo (He et al., 2020), and Barlow-twins (Zbontar et al., 2021). These methods have been used to directly apply data augmentation as training objectives (Pizzi et al., 2022; Wang et al., 2021) to address severe transformations.
ICD involves two key challenges. The first is the large-scale detecting systems. In large-scale systems, millions of images in the database are pre-processed offline into descriptors and stored. Online query images are then converted into descriptors in real-time, and a nearest-neighbor search method is employed, similar to image search systems(Liu et al., 2007; Guo et al., 2020; Johnson et al., 2019). To address this issue, lightweight architectures have been introduced to convert images into descriptors in real-time. However, several studies (Shi et al., 2022; Gu et al., 2021) have found that lightweight networks with limited representation power cannot directly employ SSL to produce an acceptable performance. Previous studies have introduced SimCLR to train ResNet-50(RN-50) and ResNeXt-101, and we find that directly applying SimCLR to lightweight networks results in poor performance. Consequently, creating smaller descriptor sizes is an effective method to reduce the storage space for descriptors. Previous method (Pizzi et al., 2022) uses principal component analysis (PCA) to create small-sized descriptors, but it leads to significant performance degradation (Pizzi et al., 2022).
A second challenge for ICD stems from the difficulty of distinguishing hard negative samples. In ICD, there are numerous hard negative samples that are visually similar to the original image but are not edited copies. These include images taken from different camera angles or images of the same location captured at different times, which can significantly complicate the identification process. In ICD, each individual instance is treated as a separate category, making it crucial for the embedding space to achieve a uniform distribution. This arrangement helps to further each instance from its nearest neighbor (hard negative). In self-supervised copy detection (SSCD) (Pizzi et al., 2022), the Kozachenko-Leononenko estimator (Sablayrolles et al., 2018) is an entropy regularizer used to promote a uniform embedding distribution, thus ensuring that the distances from embedding regions are more comparable. By achieving a uniform distribution, the approach makes full use of all dimensions, thereby addressing the issue of dimensional collapse. Through experiments, we confirmed that training a student model with a teacher model previously trained using the entropy regularizer partially alleviated the issue of dimensional collapse. However, the student model still does not fully utilize all dimensions of the descriptors, continuing to struggle with dimensional collapse.
In this work, we introduce Relational self-supervised Distillation with Compact Descriptors (RDCD), which utilizes both relation-based self-supervised distillation (RSD) and the HN loss. RSD is instrumental in accurately capturing the intricate relationships between teacher model descriptors, to ensure their effective representation in a reduced feature space. Existing SSL with knowledge distillation (SSL-KD) methods use feature-based knowledge distillation (FKD) (Gao et al., 2021) or RSD (Fang et al., [n. d.]; Abbasi Koohpayegani et al., 2020) for pre-training in classification tasks. FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error (MSE) loss. However, we found that FKD fails to prevent dimensional collapse, leading to performance degradation in ICD. We create a separate instance queue for the teacher to make the student mimic the pairwise similarity generated by the teacher through Kullback-Leibler (KL) divergence, allowing the student network to generate discriminative descriptors in its own feature space.
To reduce the descriptor size, we use a linear projector to create descriptors and conduct SSL with MoCo-v2 (Chen et al., 2020a). It creates a multilayer perceptron (MLP) by adding two linear projectors: one to enable RSD with embeddings and the other for small descriptor sizes. According to (Jing et al., 2021), using an MLP in the last layer leads to dimensional collapse due to the multiplication of the weight matrices in each linear layer, which severely degrades overall performance. We use HN loss to avoid this collapse without impairing the performance of RSD from the teacher network, thus increasing the overall performance. The HN loss function is applied to the sample with the highest similarity within the instance queue used in MoCo-v2, to decrease the similarity of the most challenging negative samples. This function encourages the network to create more discriminative descriptors between positive and negative samples, leading to a more effective feature space where dissimilar samples are further apart. We consequently achieve a comparable performance using compact descriptors with our method as shown in Figure 1.
We conduct comparative experiments with SimCLR (Chen et al., 2020b), MoCo-v2 (Chen et al., 2020a), FKD, and RSD to analyze the effectiveness of each method in training efficient lightweight networks for ICD. We test our approach on various teacher and student networks, including convolutional neural networks (CNNs) and vision transformers (ViTs), to demonstrate its architecture-agnostic characteristics. The main contributions of this work are summarized as follows:
  • We introduce RDCD, a novel approach that combines RSD with HN loss to train lightweight networks for ICD in a self-supervised manner. Our method leverages the relational information between descriptors generated by a teacher network to guide the learning of a student network with a compact descriptor size.
  • Our RSD approach effectively maintains discriminative power when mapping to a reduced feature space, overcoming the limitations encountered in FKD.
  • We employ HN loss to prevent the dimensional collapse that can occur in SSL with compact descriptors, ensuring a structured and informative feature space.
  • Through extensive experiments, we validate the effectiveness of our RDCD approach with various architectures, including CNNs and ViTs.
2. Related Work
2.1. Knowledge Distillation
Using a compact model to approximate a function learned by a more comprehensive, higher-performing model was first introduced in (Buciluǎ et al., 2006). This concept was expanded in (Hinton et al., 2015), where the student model was trained to mimic the teacher’s softened logits, a process referred to as knowledge distillation. Feature distillation has been garnering significant attention recently. (Romero et al., 2014) added a new dimension by proposing a hint-based training scheme for the alignment of feature maps, known as feature distillation, and chose the L2 distance as the metric for comparing the two feature maps. Similarly, (Zagoruyko and Komodakis, 2016) suggested the transfer of spatial attention maps from a high-performing teacher network to a smaller student network. (Yim et al., 2017) devised the idea of transferring ’flow’—defined as the inner product between features from two layers— to the student rather than knowledge. (Passalis and Tefas, 2018) directly aligned the probability distributions of the data between the teacher’s and student’s feature spaces. The success of these knowledge distillation methods can primarily be attributed to the insightful knowledge embedded in the logits of the teacher model.
2.2. Relation-based Knowledge Distillation
While conventional knowledge distillation only extracts the information for a single data point from the teacher, similarity-based distillation methods (Tung and Mori, 2019; Abbasi Koohpayegani et al., 2020; Fang et al., [n. d.]; Passalis and Tefas, 2018; Park et al., 2019; Tian et al., 2019; Tejankar et al., 2021) learn the knowledge from the teacher in terms of similarities between data points. A pairwise similarity matrix has been proposed to retain the interrelationships of similar samples between the representation space of the teacher and student models. CompRess (Abbasi Koohpayegani et al., 2020) and SEED (Fang et al., [n. d.]) which are approaches that are most closely related to our method, use similarity-based distillation to compress a large self-supervised model into a smaller one.
2.3. Self-supervised Learning
In recent studies on SSL, contrastive-based approaches have produced remarkable results in downstream tasks. Many of these techniques are derived from noise-contrastive estimators (Gutmann and Hyvärinen, 2010), where the latent distribution is estimated based on the contrast with randomly or artificially generated noises. InfoNCE (Oord et al., 2018a) was the first to propose learning image representations by predicting the future using an auto-regressive model for unsupervised learning. Subsequent studies have focused on improving the efficiency (Henaff, 2020) and utilizing multi-view as positive samples (Tian et al., 2020). However, these approaches have limited access to negative instances, because they can only rely on a limited set of negative samples. To address this limitation, (Wu et al., 2018) designed a memory queue to store previously encountered random representations as negative samples, treating each of them as independent categories, for instance discrimination. However, this approach introduces an inefficiency in which the previously stored vectors are inconsistent with the recently computed representations during the earlier stages of pre-training. To mitigate this issue, SimCLR (Chen et al., 2020b) samples negative instances from a large batch while MoCo (He et al., 2020) improves on the memory-queue-based method by utilizing a momentum-updated encoder to reduce representation inconsistency.
Figure 2. Overall pipeline of proposed Relational Self-supervised Distillation for Image Copy Detection (RDCD).
3. Methodology
3.1. Overall Architecture
The overall architecture of the proposed RDCD approach is presented in Fig 2. In RDCD, the student network is trained using three different objectives. Because the student network $f^{S}(\cdot)$ is the encoder that we want to improve, we freeze teacher network $f^{T}(\cdot)$ which is a pre-trained encoder by training using an off-the-shelf method.
Given an image x, the teacher network extracts its representation $h^{T}$, in which $h^{T}\in\mathbb{R}^{D_{T}}$ is the final feature without any classifier. The student network extracts two representations $h^{S}$ and $h^{S^{\prime}}$ from input images augmented in different ways. An FC layer is employed immediately after the student network to match the dimensionalities of $h^{T}$ and $h^{S}$. Furthermore, we also add an additional projector for contrastive learning by the student. This projector decreases the dimensions of representation and is used to calculate pairwise similarities for the computation of the contrastive loss function. We use SimCLR and MoCo-v2 as our contrastive learning methods, but any contrastive-based learning method can be adopted.
3.2. Relational Self-supervised Distillation
Relational Self-supervised Distillation (RSD) is a method that transfers instance relations from a teacher network to a student network. This process involves the creation of an instance queue within the teacher’s network, designed to store the teacher’s instance embeddings. This queue serves as a reference point for the student network, facilitating the transfer of knowledge. When a new sample is introduced, its similarity scores are calculated against all instances in the queue ($q_{k}$) using both the teacher and student networks. The similarity scores are typically computed using cosine similarity.
The key objective is to align the similarity score distribution generated by the student network with that of the teacher network. This alignment is achieved by minimizing the KL divergence between the similarity score distributions of the two networks, ensuring a close match in their respective interpretations of the instance similarities. We apply the same augmentations into an image $x_{i}$ as a batch and map the embeddings $h^{T}=f^{T}_{\theta}(\tilde{x_{i}})$ and $h^{S}=f^{S}_{\theta}(\tilde{x_{i}})$ where $h^{T},h^{S}\in\mathbb{R}^{D}$ and $f^{T}_{\theta}$ and $f^{S}_{\theta}$ denote the teacher and student network, respectively.
We compute the cosine similarity between l2-normalized descriptors from the teacher network and the queue. For the teacher and student, the similarities are:
(1) $sim(h^{T}_{i},q^{T}_{j})=\begin{bmatrix}h^{T}_{i}/\lVert h^{T}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$
(2) $sim(h^{S}_{i},q^{T}_{j})=\begin{bmatrix}h^{S}_{i}/\lVert h^{S}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$
where $sim$ represents the cosine similarity, and $q_{j}^{\top}$ denotes the transposed j-th component in the teacher’s queue $Q^{T}$ to compute cosine similarity. We calculate the probability of the i-th instance with the j-th component in the teacher’s queue.
(3) $p^{T}_{i,j}=\frac{exp(sim(h^{T}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}exp(sim(h^{T}_{i},q^{T})/\tau^{T})},\quad where\;j\in[1\dots K]$
(4) $p^{S}_{i,j}=\frac{exp(sim(h^{S}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q}exp(sim(h^{S}_{i},q^{T})/\tau^{S})},\quad where\;j\in[1\dots K]$
where $\tau^{T}$ and $\tau^{S}$ are the temperature parameters of the teacher and student networks, respectively. K is the size of the teacher’s queue. Also, $p^{T}_{i,j}$ denotes the similarity score between the embeddings $h^{T}_{i}$ from the teacher network and the embeddings in the teacher’s queue. In the same way, $p^{S}_{i,j}$ denotes the similarity score between the embeddings $h^{S}_{i}$ from the student network and the embeddings in the teacher’s queue. The objective of our RDCD is to minimize the KL-divergence between the probabilities over all input instances for the teacher and student networks. The final loss function is determined as follows:
$\mathcal{L}_{rel}$ $=\sum_{i}^{N}KL(p^{T}_{i}\ ||\ p^{S}_{i})$
$=\arg\min_{\theta_{S}}\sum_{i}^{N}\sum_{j}^{K}-\frac{\exp(\text{sim}(h^{t}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}\exp(\text{sim}(h^{t}_{i},q^{T})/\tau^{T})}\cdot$
(5) $\qquad\log\left(\frac{\exp(\text{sim}(h^{s}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q^{T}}\exp(\text{sim}(h^{s}_{i},q^{T})/\tau^{S})}\right)$
3.3. Contrastive Learning for ICD
The size of the embeddings within the architecture of lightweight networks is generally small, and it is crucial to maintain this size during training while benefiting from knowledge distillation. To ensure this, we attach a linear projector to the end of the student network that is designed to reduce the representation size, resulting in smaller embeddings. We use these final embeddings as our descriptors, which allows for simultaneous training using our relational distillation approach. We follow the MoCo-v2 (Chen et al., 2020a) training procedure by utilizing a negative instance queue generated by the momentum encoder. We employ the InfoNCE (Oord et al., 2018b) loss to guide contrastive learning, which is calculated as follows:
(6) $\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z^{S}_{i},z^{S^{\prime}}_{i})/\tau)}{\sum_{q^{S}\sim Q^{S}}\exp(\text{sim}(z^{S}_{i},q^{S})/\tau)}$
For each pair index $i$, $z^{S}_{i}$, and $z^{S^{\prime}}_{i}$ are the representations of the two augmented views of the same image (i.e., a positive pair). $\text{sim}(z^{S}_{i},z^{S^{\prime}}_{i})$ is a function measuring the similarity between these descriptors, typically a dot product. $\tau$ is a temperature parameter that scales the similarity scores. The denominator sums the exponentiated similarities of $z^{S}_{i}$ with all negative descriptors $q^{S}$ in all students’ queue $Q^{S}$.
3.4. The Hard Negative Loss
In the copy detection scenario, the primary challenges for training arise from the hard negative pairs and the fact that the use of small descriptors renders the environment susceptible to dimensional collapse. To address this, we employ the Hard Negative (HN) loss (Kordopatis-Zilos et al., 2023) to finely control the entropy and minimize the hard negative pairs. We also introduce an additional term that ensures that a hard negative pair can push their descriptors apart.
(7) $\mathcal{L}_{hn}=-\frac{1}{N}\log\sum_{i=1}^{N}max_{j\in n_{i,j}}(1-S_{i,j})$
$n_{i,j}$ denotes the row indices of the negative pairs for the i-th row of the similarity matrix.
The objective function of the student network is a combination of the contrastive loss (con), the Relational Self-supervised Distillation (rel) loss, and the hard negative (HN) loss. The final objective can be expressed as:
(8) $\mathcal{L}_{RDCD}=\lambda_{rel}\mathcal{L}_{rel}+\lambda_{con}\mathcal{L}_{con}+\lambda_{hn}\mathcal{L}_{hn}$
4. Experiments
4.1. Dataset
DISC2021 is a dataset consisting of training, reference and query images, that is used for the Image Similarity Challenge (Douze et al., 2021b). The query image set contains strong auto-augmented and human-augmented images, and distractive images that may not appear in the reference set. The training set is designed to train a model without any labels, and the reference set is used for searching the background as a database.
Copydays (Jegou et al., 2008) is a dataset designed to detect copied content, that consists of 157 original images and 2,000 queries with cropping, jpeg compressing and strong augmentations. We add 10k distractors from YFCC100M (Thomee et al., 2016), a common practice (Caron et al., 2021; Berman et al., 2019) that is known as CD10K. We evaluate our method based on the mean average precision(mAP) and micro average precision($\mu AP$) for the strongly transformed copies.
NDEC (Wang et al., 2023) dataset incorporates HN distractors, making it more advanced than conventional ICD datasets. The basic data is derived from DISC2021 while the HN images are sourced from OpenImage. Similar to DISC2021, the NDEC dataset is organized into training, query, and reference sets. The training data includes 900,000 basic images and adds HN pairs (100,000 x 2 paired images) to enhance the difficulty level of ICD. The test data contains 49,252 query images and 1,000,000 reference images, with 24,252 of the query images being HN.
4.2. Evaluation Metrics
To enable a fair comparison with state-of-the-art methods (Pizzi et al., 2022; Caron et al., 2021), we use $\mu AP$ (Perronnin et al., 2009), accuracy at 1, and recall at precision 90 as the evaluation metrics. The equation for $\mu AP$ is as follows:
(9) $\mu AP=\sum\limits_{i=1}^{N}P(i)\triangle r(i)\in[0,1]$
where $P(i)$ is the precision at position $i$ of the sorted list of pairs, $\triangle r(i)$ is the difference in the recall between position i and i-1, and N is the total number of returned pairs for all queries. Any detected pair for a distractor query will decrease the average precision, thus all queries are evaluated together by merging the returned pairs for all queries, sorting them by confidence, and generating a single precision-recall curve. This is different from mAP, also known as macro-AP (Perronnin et al., 2009), where the average precision is computed separately per query and then averaged over all queries. $\mu AP$ considers confidence values to capture all queries, while mAP only considers the query in the ground truth. Thus, $\mu AP$ is a more accurate metric in our context because our queries contain the images not contained in the references. The former is more appropriate for ICD, while the latter is typically used in retrieval tasks.
[…]
Method Network Size $\mu$AP $\mu$APSN
Multigrain RN-50 1500 16.5 36.5
Multigrain RN-50 2048 20.5 41.7
DINO ViT-B/8 1536 32.6 54.4
DINO ViT-B/16 1536 32.2 53.8
SSCD RN-50 512 61.5 72.5
SSCD ResNeXt-101 1024 63.7 75.3
using lightweight
SSCD EN-B0 64 38.2 48.5
SSCD EN-B0 128 43.5 56.2
SSCD EN-B0 256 46.0 59.8
SSCD EN-B0 512 43.6 61.1
ours
RDCD EN-B0 64 43.9 53.5
RDCD EN-B0 128 50.0 61.1
RDCD EN-B0 256 52.7 65.7
4.3. Training Implementation
We train all networks using the pretraining parameters from ImageNet-1K. Batch normalization statistics are synchronized across all GPUs. We adhere to the hyperparameters outlined in SimCLR (Chen et al., 2020b), MoCo-v2 (Chen et al., 2020a) and SSCD (Pizzi et al., 2022). The batch size is set at N = 256, with a resolution of 224 x 224 for training and 288 x 288 for inference. We use a learning rate of 1.0 and a weight decay of $10^{-6}$. We train all models for 100 epochs. We employ a cosine learning rate scheduler with a warmup period of five epochs and use the Adam optimizer. For our knowledge distillation process, we set the temperature at 0.04 for the teacher and 0.07 for the student. All experiments are conducted on an NVIDIA A100 80GB GPU. To ensure reproducibility, we maintain consistent seed values throughout all of the experiments.
4.4. Results
Table 1 presents the results for the DISC2021 dataset. The first section of the table summarizes the performance of different methods with varying descriptor sizes, while the second section describes the performance of our baseline SSCD (Pizzi et al., 2022) strategy using a lightweight network with descriptor sizes of 64, 128, and 256. When training using a lightweight network, we do not use a mix-up for the fair comparison of our methods. The third section shows the results of the proposed RDCD approach.
When compared with the methods outlined in the first section, our RDCD method achieves competitive performance despite using significantly smaller descriptors and a more compact network size. It is noteworthy that the RDCD method, with a descriptor size of 64, achieves a $\mu$APSN of 53.5, which is comparable to the DINO method that utilizes a ViT-B/16 network with a descriptor size of 1536, yielding a $\mu$APSN of 53.8. This comparison is significant, as the descriptor size of RDCD is 24 times smaller.
When evaluated against an equivalent lightweight architecture, RDCD consistently outperforms SSCD across all three dimensions assessed. We employ the SSCD model RN-50, which has a $\mu$APSN of 72.5, as the teacher network to provide knowledge to the lightweight network. When we apply our methods to EfficientNet-B0, RDCD outperforms SSCD. Remarkably, RDCD with a descriptor size of 128 achieves a $\mu$APSN of 61.1, which is competitive and matches the performance of SSCD with a descriptor size of 512. This indicates that our method is capable of delivering strong results even with smaller descriptors.
RDCD also outperforms SSCD in terms of mAP using the CD10K dataset (Table 2). Specifically, when using EfficientNet-B0 with a descriptor size of 128, RDCD achieves an mAP of 79.2, which is significantly higher than the best result for SSCD. With a descriptor size of 256, the performance improvement is even more pronounced, with RDCD reaching a mAP of 81.4.
For the NDEC dataset, Table 3 shows that our method shows competitive performance compared to the teacher network SSCD-RN50, which achieves a $\mu$APSN of 23.3 with a descriptor size of 512. Our proposed RDCD approach, employing EN-B0 network with descriptor sizes of 64, 128, and 256 achieves a $\mu$APSN of 17.3, 19.5, and 21.3, respectively. This demonstrates that across all compact descriptor sizes, RDCD surpasses the performance of the previous SSCD when using the same lightweight architecture.
Table 2. Results on the CD10K (Copydays+ 10k distractors). We compare different methods, network architectures (ResNet-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (mean Average Precision (mAP) and micro Average Precision ($\mu$AP). * means our implementation.
Method Network Size mAP $\mu$AP
Multigrain RN-50 1500 82.3 77.3
DINO ViT-B/8 1536 85.3 91.7
DINO ViT-B/16 1536 80.7 88.7
SSCD RN-50 512 85.0 97.9
SSCD ResNext-101 1024 91.9 96.5
using lightweight
SSCD EN-B0 64 64.1 95.6
SSCD EN-B0 128 71.9 97.0
SSCD EN-B0 256 76.6 97.4
SSCD EN-B0 512 77.4 97.4
ours
RDCD EN-B0 64 72.4 94.5
RDCD EN-B0 128 79.2 96.1
RDCD EN-B0 256 81.4 95.4
Table 3. Comparison of NDEC results. * means our implementation.
Method Network Size $\mu$AP $\mu$APSN
DINO ViT-B/8 1536 0.08 11.4
DINO ViT-B/16 1536 0.09 13.1
SSCD RN-50 512 21.2 23.3
using lightweight
SSCD EN-B0 64 14.4 16.4
SSCD EN-B0 128 16.5 18.7
SSCD EN-B0 256 17.5 20.1
SSCD EN-B0 512 17.3 20.4
ours
RDCD EN-B0 64 15.7 17.3
RDCD EN-B0 128 17.5 19.5
RDCD EN-B0 256 18.8 21.3
5. Discussion
5.1. Effect of RSD and HN Loss
Our method employs an additional projector to reduce the descriptor size, which creates a network with a two-layer MLP. This configuration can lead to implicit regularization, due to the interaction between the weight matrices across different layers. Implicit regularization typically restricts the ability of a network to learn diverse features, leading to dimensional collapse with contrastive SSL (Jing et al., 2021). However, when applying HN loss together with contrastive learning, we show that dimensional collapse does not occur even with the use of an MLP, thus reducing the impact of implicit regularization.
Figure 3. Log of singular values for a descriptor size of 256, with and without HN loss.
To verify this, we calculate the singular values of the final descriptor of a model with a descriptor size of 256 under four conditions: with the use of RSD, FKD and with and without the use of HN loss. We compute the descriptors for 50k queries from the DISC21 dataset, calculate their covariance, and perform singular value decomposition on the covariance matrix. Subsequently, we extract the singular values and apply a logarithmic operation to them. As illustrated in Figure 3, RDCD has a full rank, while the absence of HN loss results in dimensional collapse. It can also be observed that applying HN loss with FKD does not have a significant effect.
During the training, we visualize the similarity of the hard negative for both the MoCo-v2 baseline and HN loss, implemented with either RSD or FKD, using descriptor sizes of 128 or 256, as shown in Figure 4. FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error (MSE) loss. As Figure 4 shows, using FKD with HN loss fails to reduce the similarity of the nearest negatives, whereas using RSD with HN loss gradually reduces the nearest negative similarity.
We further visualize the difference between positive similarity and hard negative similarity across multiple dimensions experimented in MoCo-v2 — RSD environment, comparing scenarios with and without the application of HN loss. We analyze a sample of 5,000 distinct queries extracted from the DISC21 dataset, which are separate from the queries used in the standard evaluation process for generality, to calculate and assess the differences in the similarity between positive and hard negative samples. As Figure 5 shows, across all descriptor sizes, the application of HN loss resulted in a more pronounced disparity compared to when it was omitted. This observation implies that HN loss effectively enhances the separation between positive and hard negative samples. Furthermore, we noted that as the dimensionality increases, the difference between positive and hard negative similarity also enlarges. We conjecture that this trend may contribute to improved performance in higher-dimensional descriptors.
Figure 4. Comparison of hard negative similarity of RSD and FKD.
(a) 64 descriptors
(b) 128 descriptors
(c) 256 descriptors
Figure 5. Comparison of the difference in similarity between positives and nearest negatives with and without the use of HN Loss.
Table 4. Comparison of other distillation. All methods are trained with EfficientNet-B0 network. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
128 descriptors
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD 43.5 56.2 - -
SimCLR 46.2 58.7 48.4 59.9
47.4 59.6 49.6 61.4
47.5 59.6 50.8 61.5
MoCo-v2 40.3 57.5 44.8 59.1
47.1 60.7 50.0 61.1
46.2 59.3 49.6 61.0
256 descriptors
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD 46.0 59.8 - -
SimCLR 51.7 64.0 51.4 64.5
52.2 65.6 52.7 65.8
52.4 65.5 49.9 64.6
MoCo-v2 41.3 56.6 46.1 60.1
47.2 61.0 52.9 65.7
47.3 60.8 53.2 65.9
5.2. Comparison with Other Forms of Distillation
To further demonstrate the efficacy of RSD, we conduct a comparative analysis with FKD by 1) employing FKD alone, 2) employing RSD alone, and 3) integrating both FKD and RSD (Table 13). We also employ SimCLR-style contrastive learning. When FKD and RSD are combined, the best performance is observed in several cases, suggesting that combining both distillation methods can further improve performance. Nevertheless, RSD consistently outperforms FKD in all cases, with RSD alone achieving performance on par with the combined use of both distillation methods. This indicates that RSD is sufficiently robust, and FKD does not significantly alter the outcome. It confirms that RSD can operate alone to enhance the performance of compact descriptors. Additional experiments with other distillation methods are presented in Appendix A.
5.3. Effect of Different Architectures
We demonstrate the effectiveness of our proposed approach in scenarios where the teacher networks belong to different architectural families (ViT or CNN), with student models also selected from these two distinct architectures. For the ViT, we use DINO as the teacher, and for the CNNs, we use SSCD-RN50 and SSCD-ResNeXt-101. All models are trained on SSCD-RN50 hyperparameter settings. With DINO as the teacher model, we follow the ICD settings presented in (Caron et al., 2021) for the DINO baseline. Consequently, we distill a descriptor size of 1536 from the teacher (concatenation of 768 class tokens and 768 GeM pooled patch tokens). In the case of ResNeXt-101 teacher model, we distill the final target descriptor with a size of 1024. Table 5 shows that our method is highly effective for the different architectures, highlighting its adaptability and robustness.
Table 5. DISC evaluation with different architectures. All experiments are performed with a descriptor size of 128.
T S Eff-B0 FastViT-T12 Mob-V3
$\mu$AP $\mu$APSN $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
DINO-ViT-B/8 32.1 52.1 28.1 52.2 39.6 54.7
SSCD-RN-50 50.0 61.1 48.5 61.1 47.3 58.8
SSCD-ResNeXt-101 44.9 59.9 42.8 57.4 41.1 56.6
5.4. Ablation Study
Effectiveness of Each Loss Term. To further assess the effectiveness of RDCD, we investigate the impact of different loss components on its performance, specifically, RSD loss and HN loss as shown in Table 6. In this section, we present an additional metric, the rank preserving ratio (RPR), which is the rank divided by the descriptor dimension. A low RPR indicates dimensional collapse, and RPR value of 1 indicates that the descriptor makes full use of all dimensions. All models are trained on the DISC21 dataset with 100 epochs and SSCD-RN50-w/mixup is employed as the teacher model. Our result shows that with the addition of RSD loss and HN loss, the performance improves substantially, which indicates that RSD and HN loss can benefit the performance of RDCD.
Table 6. Ablation study on the loss terms in RDCD. Descriptor sizes of 64, 128, and 256 were examined, and a consistent MLP configuration was maintained across all experiments for a fair comparison.
Loss MLP-d $\mu$AP $\mu$APSN
$\mathcal{L}_{con}$ $\mathcal{L}_{rel}$ $\mathcal{L}_{hn}$
MoCo-v2
1280/512/64 10.9 14.4
1280/512/128 11.9 17.2
1280/512/256 16.5 26.4
Effectiveness of RSD loss
1280/512/64 39.9 52.1
1280/512/128 47.1 60.7
1280/512/256 47.2 61.0
Effectiveness of HN loss
1280/512/64 43.9 53.5
1280/512/128 50.0 61.1
1280/512/256 52.9 65.7
Influence of $\lambda_{rel}$ and $\lambda_{hn}$. $\lambda_{rel}$ and $\lambda_{hn}$ are the weights of RSD loss and HN loss, respectively. First, we analyze the influence of contrastive loss and RSD loss (Table 7). As shown in the third row, a contrastive loss of 1 and RSD loss of 10 achieves the best performance, which indicates that knowledge distillation is essential for the model with better performance. The HN loss is also varied with the contrastive loss and RSD loss fixed at 1 and 10, respectively, showing that the best performance occurs when the HN loss is 5. However, for larger values, the performance drops to 0, (i.e. collapses) because the network focuses more on HN loss rather than contrastive and distillation loss. We further calculate RPR of each experiment. In the first section, where only contrastive loss and RSD loss are employed, we observe that RPR gradually increases incrementally with an increase in $\lambda_{rel}$. This suggests that the student model partially alleviates the dimensional collapse which is guided by a teacher model that has been trained with an entropy regularizer (SSCD). Subsequently, in the second section, where HN loss is incorporated, the student model finally utilizes the full dimensions of the descriptors.
Table 7. Ablation study to evaluate the impact of the loss ratio. All experiments are performed with a descriptor size of 256.
$\lambda_{con}$ $\lambda_{rel}$ $\lambda_{hn}$ RPR $\mu$AP $\mu$APSN
10 1 - 0.27 17.0 32.5
1 1 - 0.40 33.4 52.4
1 10 - 0.55 47.2 61.0
1 10 1 0.62 47.3 61.4
1 10 3 0.98 47.1 62.1
1 10 5 1.0 52.9 65.7
1 10 above 5 - collapse collapse
We next conduct a step-by-step experiment with DINO as the teacher model (Table 8). Our approach reduces the descriptor size by a factor of 12 but produces a performance comparable to that of the teacher model. As illustrated in Discussion 5.1, dimensional collapse occurs when HN loss is not employed. Applying HN loss thus successfully improves performance and prevents dimensional collapse.
Table 8. Ablation study of using DINO as teacher model.
Method Network Size RPR $\mu$AP $\mu$APSN
DINO (Teacher) ViT-B/8 1536 0.99 32.6 54.4
RSD EN-B0 1536 0.88 10.0 35.2
RSD — MoCo-v2 EN-B0 128 0.57 20.7 40.0
RSD — MoCo-v2 — HN (RDCD) EN-B0 128 1.0 32.1 52.1
6. Conclusion
In this paper, we present RDCD, a novel method for training lightweight networks with compact descriptors for image copy detection in a self-supervised manner. We demonstrate in a series of experiments that RDCD, which combines RSD with HN loss effectively prevents dimensional collapse in lightweight architectures, and achieves a competitive performance across various benchmarks. We believe that this approach offers significant advantages in search speed and scalability for multimedia applications.
Appendix
Appendix A Different Architectures
To show the generality of our method, we adopt one lightweight model each from the most commonly used architectures in computer vision, CNN and ViT, conducting experiments with MobileNet-V3 and FastViT-T12. Our results are shown in Table 9, Table 10, Table 11.
Table 9. Performance of DISC2021 with different student architectures. * means our implementation.
Method Network Size $\mu$AP $\mu$APSN
using lightweight
SSCD Mob-V3 128 45.5 56.2
SSCD FastViT-T12 128 42.5 59.7
SSCD Mob-V3 256 47.3 60.1
SSCD FastViT-T12 256 43.1 59.5
ours
RDCD Mob-V3 128 50.6 60.8
RDCD FastViT-T12 128 48.5 61.1
RDCD Mob-V3 256 53.9 65.6
RDCD FastViT-T12 256 56.4 67.4
Table 10. Performance of CD10K(Copydays+ 10k distractors) with different student architectures. * means our implementation.
Method Network Size mAP $\mu$AP
Multigrain RN-50 1500 82.3 77.3
DINO ViT-B/8 1536 85.3 91.7
DINO ViT-B/16 1536 80.7 88.7
SSCD RN-50 512 85.0 97.9
SSCD ResNext-101 1024 91.9 96.5
using lightweight
SSCD Mob-V3 128 68.9 97.1
SSCD FastViT-T12 128 60.9 87.0
SSCD Mob-V3 256 75.4 97.4
SSCD FastViT-T12 256 74.4 92.2
ours
RDCD Mob-V3 128 81.3 96.5
RDCD FastViT-T12 128 77.4 95.3
RDCD Mob-V3 256 83.5 97.6
RDCD FastViT-T12 256 79.3 97.0
Table 11. Performance of NDEC with different student architectures. * means our implementation.
Method Network Size $\mu$AP $\mu$APSN
DINO ViT-B/8 1536 16.2 22.8
DINO ViT-B/16 1536 18.1 26.2
SSCD RN-50 512 42.4 46.6
using lightweight
SSCD Mob-V3 128 34.6 38.9
SSCD FastViT-T12 128 33.5 38.6
SSCD Mob-V3 256 36.8 41.5
SSCD FastViT-T12 256 33.3 38.3
ours
RDCD Mob-V3 128 35.5 39.3
RDCD FastViT-T12 128 35.1 38.7
RDCD Mob-V3 256 37.9 42.6
RDCD FastViT-T12 256 39.1 43.2
Appendix B Comparison with Barlow-Distill
We introduce the RDCD method using the MoCo framework’s negative sample queue in our distillation process. Alternatively, there are self-supervised learning (SSL) methods that do not employ negative samples and instead focus solely on positive samples. Inspired by the Barlow Twins approach, we have integrated the capability of distillation in our process by generating positive pairs within a batch through augmentation and optimizing the cross-correlation matrix between these pairs. This method, we call Barlow-Distill, ensures that each feature operates as independently as possible from the others, thereby minimizing feature redundancy. This minimization ensures that each feature retains unique information, leading to a richer data representation.
The Barlow-Distill approach distinctly avoids the use of negative samples, with all learning occurring through the adjustment of the correlation between positive pairs. This method is particularly effective when generating different transformations of the same image through augmentation techniques. Consequently, the model enhances the representation of the same object or scene in various forms. Barlow-Distill generalizes well by minimizing local features in two distinct augmented images. This process effectively captures higher levels of abstract information, ensuring that the latent representation retains sufficient information to learn useful representations for various downstream tasks. However, this concentration on high-level features may result in the loss of crucial low-level details, which can degrade performance. This loss is particularly problematic in image copy detection tasks, where missing fine-grained visual details can lead to degraded performance.
In contrast, SSL methods utilizing contrastive learning could be more suitable for tasks such as image copy detection. This method, by clearly differentiating between positive and negative samples, enables the model to more effectively recognize and learn local detail and fine-grained visual differences. Contrastive learning enhances the model’s ability to distinguish between similar images, thus improving accuracy and efficiency in the image copy detection task. This enhancement is vital for capturing important local features in images, a crucial ability for image copy detection. According to our research results presented in Table 12, the RDCD model shows enhanced performance compared to Barlow-Distill.
Table 12. Comparision with RDCD and Barlow-Distill
Method Network Size $\mu$AP $\mu$APSN
Barlow-Distill EN-B0 64 44.1 48.3
Barlow-Distill EN-B0 128 51.9 59.0
Barlow-Distill EN-B0 256 54.8 64.4
RDCD EN-B0 64 43.9 53.5
RDCD EN-B0 128 50.0 61.1
RDCD EN-B0 256 52.7 65.7
Appendix C Intermediate Descriptor
In this study, we evaluate the performance of DISC2021 using an intermediate descriptor designed to align the descriptor size with that of the teacher model. We assess a 512-dimensional descriptor both with and without the application of HN loss, utilizing the EfficientNet-B0 network architecture. Our results (Table 13) show that the proposed RDCD method consistently outperforms SSCD across two projection sizes, 128 and 256. Furthermore, our analysis reveals that the performance gap between using and not using HN loss is negligible, and the intermediate descriptor fully exploits its 512-dimensional capacity without HN loss. This finding suggests that the student model is effectively trained under the guidance of a teacher model equipped with an entropy regularizer.
Table 13. Comparison of descriptor size of 512. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
128 projection
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD(RN50) w/o mixup 60.4 71.1 - -
SSCD(EN-B0) w/o mixup 43.6 61.1 - -
SimCLR 56.1 68.0 55.2 67.6
56.3 69.7 56.3 69.7
56.6 69.5 56.7 69.6
MoCo-v2 51.2 67.3 54.4 67.8
55.1 69.1 55.5 69.1
55.2 69.1 55.8 68.9
256 projection
Hard Negative Loss No Yes
Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
SSCD(RN50) w/o mixup 60.4 71.1 - -
SSCD(EN-B0) w/o mixup 43.6 61.1 - -
SimCLR 56.2 68.1 55.5 67.1
56.3 69.5 56.4 69.5
55.7 67.9 54.1 68.5
MoCo-v2 53.3 67.4 54.3 67.7
53.7 68.2 55.4 68.9
55.8 69.2 56.1 69.0
Appendix D Qualitative Examples
In Figure 6, we present the queries along with the top-2 results retrieved by both the RDCD and SSCD methods for cases where both methods accurately identify the ground truth. However, it is noteworthy that RDCD consistently retrieves a smaller similarity score difference between the ground truth and the hardest negative sample, compared to SSCD. For our matching evaluation, we employ a descriptor of size 256, utilizing the EN-B0 network architecture.
Figure 6. Example of top-2 retrieval results from the DISC2021 dataset.
Appendix E PCA
The dimensionality of the descriptor plays a crucial role in optimizing the trade-off between matching time and accuracy during the matching phase of ICD. Previous approaches have employed PCA and PCA whitening to create compact descriptors. The post-processing technique of PCA whitening is known for its efficacy in making the descriptor distribution more uniform (Pizzi et al., 2022). Our RDCD method, trained via deep learning, exhibits superior performance compared to the baseline, which utilizes large-sized descriptors that have undergone PCA whitening. These results are detailed in Table  14.
Table 14. The following table demonstrates that when dimension reduction is applied to the descriptor through PCA whitening, our approach outperforms the existing baseline in terms of $\mu$AP performance. We do not apply Mixup. * means our implementation.
Method Network Descriptor Size $\mu$AP $\mu$APSN $\mu$APSN256 $\mu$APSN128 $\mu$APSN64
DINO ViT-B/8 1536 32.6 54.4 47.7 42.3 32.7
DINO ViT-B/16 1536 32.2 53.8 47.4 42.0 32.0
SSCD RN-50 512 43.6 72.5 66.0 56.3 38.3
SSCD ResNext-101 1024 63.7 75.3 65.5 54.0 34.8
using lightweight
SSCD EN-B0 64 38.2 48.5 48.5 48.5 48.5
SSCD EN-B0 128 43.5 56.2 56.2 56.2 44.4
SSCD EN-B0 256 46.0 59.8 59.8 53.8 42.3
SSCD EN-B0 512 43.6 61.1 58.0 52.0 41.9
Ours
RDCD EN-B0 64 43.9 53.5 53.5 53.5 53.5
RDCD EN-B0 128 50.0 61.1 61.1 61.1 53.1
RDCD EN-B0 256 52.7 65.7 65.7 61.5 52.2

### v5

@@
 Relational Self-supervised Distillation with Compact Descriptors for Image Copy Detection
-Abstract.
-This paper addresses image copy detection, a task in online sharing platforms for copyright protection.
-While previous approaches have performed exceptionally well, the large size of their networks and descriptors remains a significant disadvantage, complicating their practical application.
-In this paper, we propose a novel method that achieves a competitive performance by using a lightweight network and compact descriptors.
-By utilizing relational self-supervised distillation to transfer knowledge from a large network to a small network, we enable the training of lightweight networks with a small descriptor size.
-Our approach, which we call Relational self-supervised Distillation with Compact Descriptors (RDCD), introduces relational self-supervised distillation (RSD) for flexible representation in a smaller feature space and applies contrastive learning with a hard negative (HN) loss to prevent dimensional collapse.
-We demonstrate the effectiveness of our method using the DISC2021, Copydays, and NDEC benchmark datasets, with which our lightweight network with compact descriptors achieves a competitive performance.
-For the DISC2021 benchmark, ResNet-50/EfficientNet-B0 are used as a teacher and student respectively, the micro average precision improved by 5.0%/4.9%/5.9% for 64/128/256 descriptor sizes compared to the baseline method.
-Knowledge Distillation, Self-supervised Learning, Lightweight Network, Image Copy Detection
-1. Introduction
-Figure 1. Comparison of RDCD(Ours) and other image copy detection methods. RDCD utilizes a lightweight network and achieves high performance despite its compact descriptor sizes.
-Visual copy detection is a task that determines whether two pieces of multimedia items originate from a single source.
-This is widely used in online sharing platforms, including social networks, to filter content for copyright protection.
-Image copy detection (ICD), a part of visual copy detection, is the task of detecting copied images from an existing database, where each image instance is treated as an individual category such as instance matching problem (Douze et al., 2021a).
-This task has also been explored under the name of near-duplicate image detection (Chum et al., 2008) or content-based image retrieval (Kim, 2003).
-In this context, their approaches are centered on category-level recognition or instance-level recognition, focusing on identifying broad categories or specific objects, respectively.
-However, ICD addresses a more nuanced challenge by focusing on exact image copies that often involve severe transformations such as re-encoding, resizing, merging, cropping, warping, or color distortion.
-Traditional approaches in this domain have primarily employed either global descriptors (Kim, 2003; Wang and Jiang, 2015) or local descriptors such as bag-of-words (Zhou et al., 2010), which may struggle with such transformations.
-To better address these challenges, recent advancements have leveraged self-supervised learning (SSL) methods such as DINO (Caron et al., 2021), SimCLR (Chen et al., 2020b), MoCo (He et al., 2020), and Barlow-twins (Zbontar et al., 2021).
-These methods have been used to directly apply data augmentation as training objectives (Pizzi et al., 2022; Wang et al., 2021) to address severe transformations.
+Abstract
+Image copy detection is the task of detecting edited copies of any image within a reference database.
+While previous approaches have shown remarkable progress, the large size of their networks and descriptors remains a disadvantage, complicating their practical application.
+In this paper, we propose a novel method that achieves competitive performance by using a lightweight network and compact descriptors.
+By utilizing relational self-supervised distillation to transfer knowledge from a large network to a small network, we enable the training of lightweight networks with smaller descriptor sizes.
+We introduce relational self-supervised distillation for flexible representation in a smaller feature space and apply contrastive learning with a hard negative loss to prevent dimensional collapse.
+For the DISC2021 benchmark, ResNet-50 and EfficientNet-B0 are used as the teacher and student models, respectively, with micro average precision improving by 5.0%/4.9%/5.9% for 64/128/256 descriptor sizes compared to the baseline method.
+The code is available at https://github.com/juntae9926/RDCD.
+Figure 1: Comparison of RDCD(Ours) and other image copy detection methods. RDCD utilizes a lightweight network and achieves high performance with compact descriptor sizes.
+1 Introduction
+Image copy detection (ICD) is the task of identifying edited copies of images within an existing database.
+This is widely used in online sharing platforms and social networks to filter content for copyright protection.
+Over the past years, ICD has been considered a sub-task of instance-level image retrieval [8].
+These approaches utilize global descriptors [2, 33, 32, 24] and local descriptors [25, 37, 42], focusing on identifying specific objects that represent the images.
+However, ICD addresses a more complex challenge by focusing on detecting exact copies of the images themselves, rather than identifying representative objects.
+It deals with precise duplicates of images that undergo severe transformations such as re-encoding, resizing, merging, cropping, warping, or color distortion.
+To better address these challenges, recent studies have leveraged self-supervised learning (SSL) methods [5, 6, 15, 49] to directly apply data augmentation as their training objective [31, 45].
 ICD involves two key challenges.
-The first is the large-scale detecting systems.
+The first challenge is the large scale of the detection system.
 In large-scale systems, millions of images in the database are pre-processed offline into descriptors and stored.
-Online query images are then converted into descriptors in real-time, and a nearest-neighbor search method is employed, similar to image search systems(Liu et al., 2007; Guo et al., 2020; Johnson et al., 2019).
-To address this issue, lightweight architectures have been introduced to convert images into descriptors in real-time.
-However, several studies (Shi et al., 2022; Gu et al., 2021) have found that lightweight networks with limited representation power cannot directly employ SSL to produce an acceptable performance.
-Previous studies have introduced SimCLR to train ResNet-50(RN-50) and ResNeXt-101, and we find that directly applying SimCLR to lightweight networks results in poor performance.
-Consequently, creating smaller descriptor sizes is an effective method to reduce the storage space for descriptors.
-Previous method (Pizzi et al., 2022) uses principal component analysis (PCA) to create small-sized descriptors, but it leads to significant performance degradation (Pizzi et al., 2022).
-A second challenge for ICD stems from the difficulty of distinguishing hard negative samples.
-In ICD, there are numerous hard negative samples that are visually similar to the original image but are not edited copies.
-These include images taken from different camera angles or images of the same location captured at different times, which can significantly complicate the identification process.
-In ICD, each individual instance is treated as a separate category, making it crucial for the embedding space to achieve a uniform distribution.
-This arrangement helps to further each instance from its nearest neighbor (hard negative).
-In self-supervised copy detection (SSCD) (Pizzi et al., 2022), the Kozachenko-Leononenko estimator (Sablayrolles et al., 2018) is an entropy regularizer used to promote a uniform embedding distribution, thus ensuring that the distances from embedding regions are more comparable.
-By achieving a uniform distribution, the approach makes full use of all dimensions, thereby addressing the issue of dimensional collapse.
-Through experiments, we confirmed that training a student model with a teacher model previously trained using the entropy regularizer partially alleviated the issue of dimensional collapse.
-However, the student model still does not fully utilize all dimensions of the descriptors, continuing to struggle with dimensional collapse.
-In this work, we introduce Relational self-supervised Distillation with Compact Descriptors (RDCD), which utilizes both relation-based self-supervised distillation (RSD) and the HN loss.
-RSD is instrumental in accurately capturing the intricate relationships between teacher model descriptors, to ensure their effective representation in a reduced feature space.
-Existing SSL with knowledge distillation (SSL-KD) methods use feature-based knowledge distillation (FKD) (Gao et al., 2021) or RSD (Fang et al., [n.
-d.]; Abbasi Koohpayegani et al., 2020) for pre-training in classification tasks.
-FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error (MSE) loss.
+Online query images are then converted into descriptors in real-time, after which a nearest-neighbor search method is employed, similar to image search systems [22, 14, 19].
+However, this large-scale approach often leads to significant computational overhead, resulting in prolonged processing times, especially when dealing with vast image databases.
+To efficiently improve large-scale systems, an ideal approach is to use smaller architectures or utilize smaller-sized descriptors.
+This approach effectively reduces search time and storage space while maintaining system performance.
+However, several studies [36, 13] have found that lightweight networks with limited representation power cannot directly employ SSL to produce acceptable performance.
+Another method uses principal component analysis (PCA) to create small-sized descriptors, but it leads to significant performance degradation [31].
+The second major challenge is the complex task of distinguishing between genuine copies and visually similar but unrelated images, known as hard negative samples.
+These hard negative samples are prevalent in ICD scenarios and can significantly complicate the identification process.
+Examples include images captured from different camera angles or images captured at the same location at different times, which are visually similar to the original but are not actual edited copies.
+A uniform distribution helps to maximize the separation between each image and its closest neighbors, including hard negatives.
+SSCD [31] employs the Kozachenko-Leononenko estimator [35] as an entropy regularizer to promote this uniform distribution.
+This technique ensures the distances between different regions in the embedding space are more consistent and comparable, thereby improving the system’s ability to differentiate between genuine copies and hard negative samples.
+By achieving uniform distribution, this approach makes full use of all dimensions, thereby addressing the issue of dimensional collapse.
+Despite these advancements, challenges still remain in efficiently training lightweight networks for ICD, especially when dealing with hard negatives, while maintaining high performance.
+In this work, we introduce Relational Self-supervised Distillation with Compact Descriptors (RDCD), a novel approach that combines relation-based self-supervised distillation (RSD) and Hard Negative (HN) loss to address these challenges.
+Our method aims to effectively capture the intricate relationships between teacher model descriptors and represent them in a reduced feature space.
+Existing SSL with knowledge distillation (SSL-KD) approaches use feature-based knowledge distillation (FKD) [12] or RSD [10, 1] for pre-training in classification tasks.
+FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error loss.
 However, we found that FKD fails to prevent dimensional collapse, leading to performance degradation in ICD.
-We create a separate instance queue for the teacher to make the student mimic the pairwise similarity generated by the teacher through Kullback-Leibler (KL) divergence, allowing the student network to generate discriminative descriptors in its own feature space.
-To reduce the descriptor size, we use a linear projector to create descriptors and conduct SSL with MoCo-v2 (Chen et al., 2020a).
-It creates a multilayer perceptron (MLP) by adding two linear projectors: one to enable RSD with embeddings and the other for small descriptor sizes.
-According to (Jing et al., 2021), using an MLP in the last layer leads to dimensional collapse due to the multiplication of the weight matrices in each linear layer, which severely degrades overall performance.
-We use HN loss to avoid this collapse without impairing the performance of RSD from the teacher network, thus increasing the overall performance.
-The HN loss function is applied to the sample with the highest similarity within the instance queue used in MoCo-v2, to decrease the similarity of the most challenging negative samples.
-This function encourages the network to create more discriminative descriptors between positive and negative samples, leading to a more effective feature space where dissimilar samples are further apart.
-We consequently achieve a comparable performance using compact descriptors with our method as shown in Figure 1.
-We conduct comparative experiments with SimCLR (Chen et al., 2020b), MoCo-v2 (Chen et al., 2020a), FKD, and RSD to analyze the effectiveness of each method in training efficient lightweight networks for ICD.
-We test our approach on various teacher and student networks, including convolutional neural networks (CNNs) and vision transformers (ViTs), to demonstrate its architecture-agnostic characteristics.
+To address this issue, we create a separate instance queue for the teacher to make the student mimic the pairwise similarity generated by the teacher using Kullback-Leibler (KL) divergence, allowing the student network to generate discriminative descriptors in its own feature space.
+Although our distillation method is similar to previous SSL-KD methods [10, 1], it ultimately differs in the type of information it aims to distill.
+The main purpose of these methods is to provide advantages for downstream tasks as pre-training methods in lightweight networks.
+Their objective is to enrich the intermediate representations with transformation-covariant features, embedding useful information, such as color and orientation [6].
+In contrast, our objective is to learn transformation-invariant representations at the final layer rather than obtaining transformation-covariant intermediate features.
+This design is crucial as the ICD task requires representations that go beyond instance discrimination to a tighter level, which is distinctly different from instance retrieval or classification tasks.
+In the ICD task, transformation-invariant features are essential, while intermediate representations with transformation-covariant characteristics degrade performance, as illustrated in [31].
+We conduct comparative experiments with SimCLR [6], MoCo-v2 [7], FKD, and RSD to analyze the effectiveness of each method in training efficient lightweight networks for ICD.
+We test our approach on various teacher and student networks, including convolutional neural networks and vision transformers, to demonstrate its architecture-agnostic characteristics.
 The main contributions of this work are summarized as follows:
-  • We introduce RDCD, a novel approach that combines RSD with HN loss to train lightweight networks for ICD in a self-supervised manner.
+  • We introduce RDCD, a novel approach that combines RSD and HN loss to train lightweight networks for ICD in a self-supervised manner.
 Our method leverages the relational information between descriptors generated by a teacher network to guide the learning of a student network with a compact descriptor size.
-  • Our RSD approach effectively maintains discriminative power when mapping to a reduced feature space, overcoming the limitations encountered in FKD.
-  • We employ HN loss to prevent the dimensional collapse that can occur in SSL with compact descriptors, ensuring a structured and informative feature space.
+  • We employ HN loss to prevent the dimensional collapse that can occur in SSL using compact descriptors, ensuring a structured and informative feature space.
   • Through extensive experiments, we validate the effectiveness of our RDCD approach with various architectures, including CNNs and ViTs.
-2. Related Work
-2.1. Knowledge Distillation
-Using a compact model to approximate a function learned by a more comprehensive, higher-performing model was first introduced in (Buciluǎ et al., 2006).
-This concept was expanded in (Hinton et al., 2015), where the student model was trained to mimic the teacher’s softened logits, a process referred to as knowledge distillation.
+Figure 2: The overall pipeline of proposed Relational Self-supervised Distillation for Image Copy Detection (RDCD).
+The method combines three key components: (1) Relational Self-supervised Distillation (RSD) which transfers knowledge from a pre-trained teacher network to a lightweight student network, (2) Contrastive Learning which uses MoCo-v2, and (3) Hard Negative (HN) Loss to address challenging negative samples.
+The student network $f^{S}_{\theta}$ extracts features $h^{S}$ and $h^{S^{\prime}}$ from two augmented views of an input image.
+These features are then projected to a lower dimension for use in contrastive learning.
+The teacher network $f^{T}_{\theta}$ guides RSD by comparing similarities between instances in its feature space.
+The final RDCD loss is a weighted combination of RSD loss ($\mathcal{L}_{rel}$), contrastive loss ($\mathcal{L}_{con}$), and HN loss ($\mathcal{L}_{hn}$), enabling the student to learn compact yet effective descriptors for image copy detection.
+2 Related Work
+2.1 Existing Image Copy Detection Methods
+The ICD task has seen significant advancements since the 2021 Image Similarity Challenge [9] hosted by Meta AI Research, which spurred numerous research developments.
+Previous approaches utilize either contrastive self-supervised learning or deep metric learning.
+CNNCL [47] uses contrastive loss and cross-batch memory.
+EFNets [27] employs Arcface loss-based metric learning and uses a drip training technique to incrementally increase the number of classes during training.
+BOT [23] uses deep metric learning through hard sample mining with triplet loss and classification loss.
+In subsequent years, SSCD [31] applies InfoNCE loss with entropy regularization.
+Different from previous winning solutions which use numerous techniques to enhance performance, such as ensembles and post-processing, our approach follows the same evaluation dataset and basic evaluation methods from SSCD [31].
+Additionally, we trained our model using a purely unlabeled dataset, without employing any ground truth labels.
+Active Image Indexing [11] applies imperceptible modifications to images called ”activation”, when using an ICD model in conjunction with Approximate Nearest Neighborhood indexing, ensuring the transformed image is positioned at the center of the quantized cluster.
+To address efficiency, our method aims to directly reduce descriptor sizes without relying on additional indexing techniques [38, 20].
+2.2 Knowledge Distillation
+Using a compact model to approximate a function learned by a more comprehensive, higher-performing model was first introduced in [4].
+This concept was expanded in [16], where the student model is trained to mimic the teacher’s softened logits, a process known as knowledge distillation.
 Feature distillation has been garnering significant attention recently.
-(Romero et al., 2014) added a new dimension by proposing a hint-based training scheme for the alignment of feature maps, known as feature distillation, and chose the L2 distance as the metric for comparing the two feature maps.
-Similarly, (Zagoruyko and Komodakis, 2016) suggested the transfer of spatial attention maps from a high-performing teacher network to a smaller student network.
-(Yim et al., 2017) devised the idea of transferring ’flow’—defined as the inner product between features from two layers— to the student rather than knowledge.
-(Passalis and Tefas, 2018) directly aligned the probability distributions of the data between the teacher’s and student’s feature spaces.
+FitNet [34] adds a new dimension by proposing a hint-based training scheme for the alignment of feature maps, known as feature distillation, selecting the L2 distance as the metric for comparing the two feature maps.
+Similarly, [48] suggests the transfer of spatial attention maps from a high-performing teacher network to a smaller student network.
+[46] introduces the idea of transferring ’flow’—defined as the inner product between features from two layers— to the student.
+PKT [29] directly aligns the probability distributions of the data between the teacher’s and student’s feature spaces.
 The success of these knowledge distillation methods can primarily be attributed to the insightful knowledge embedded in the logits of the teacher model.
-2.2. Relation-based Knowledge Distillation
-While conventional knowledge distillation only extracts the information for a single data point from the teacher, similarity-based distillation methods (Tung and Mori, 2019; Abbasi Koohpayegani et al., 2020; Fang et al., [n.
-d.]; Passalis and Tefas, 2018; Park et al., 2019; Tian et al., 2019; Tejankar et al., 2021) learn the knowledge from the teacher in terms of similarities between data points.
+2.3 Relation-based Knowledge Distillation
+While conventional knowledge distillation only extracts the information for a single data point from the teacher, similarity-based distillation methods [43, 1, 10, 29, 28, 41, 39] learn the knowledge from the teacher in terms of similarities between data points.
 A pairwise similarity matrix has been proposed to retain the interrelationships of similar samples between the representation space of the teacher and student models.
-CompRess (Abbasi Koohpayegani et al., 2020) and SEED (Fang et al., [n.
-d.]) which are approaches that are most closely related to our method, use similarity-based distillation to compress a large self-supervised model into a smaller one.
-2.3. Self-supervised Learning
-In recent studies on SSL, contrastive-based approaches have produced remarkable results in downstream tasks.
-Many of these techniques are derived from noise-contrastive estimators (Gutmann and Hyvärinen, 2010), where the latent distribution is estimated based on the contrast with randomly or artificially generated noises.
-InfoNCE (Oord et al., 2018a) was the first to propose learning image representations by predicting the future using an auto-regressive model for unsupervised learning.
-Subsequent studies have focused on improving the efficiency (Henaff, 2020) and utilizing multi-view as positive samples (Tian et al., 2020).
-However, these approaches have limited access to negative instances, because they can only rely on a limited set of negative samples.
-To address this limitation, (Wu et al., 2018) designed a memory queue to store previously encountered random representations as negative samples, treating each of them as independent categories, for instance discrimination.
-However, this approach introduces an inefficiency in which the previously stored vectors are inconsistent with the recently computed representations during the earlier stages of pre-training.
-To mitigate this issue, SimCLR (Chen et al., 2020b) samples negative instances from a large batch while MoCo (He et al., 2020) improves on the memory-queue-based method by utilizing a momentum-updated encoder to reduce representation inconsistency.
-Figure 2. Overall pipeline of proposed Relational Self-supervised Distillation for Image Copy Detection (RDCD).
-3. Methodology
-3.1. Overall Architecture
+CompRess [1] and SEED [10], which are approaches related to our method, use similarity-based distillation to compress a large self-supervised model into a smaller one.
+However, our method differs in that we distill the relationship at the final layer to learn the transformation-invariant descriptor itself, whereas previous approaches enrich the intermediate representations with more transformation-covariant features.
+3 Methodology
+3.1 Overall Architecture
 The overall architecture of the proposed RDCD approach is presented in Fig 2.
 In RDCD, the student network is trained using three different objectives.
-Because the student network $f^{S}(\cdot)$ is the encoder that we want to improve, we freeze teacher network $f^{T}(\cdot)$ which is a pre-trained encoder by training using an off-the-shelf method.
+Because the student network $f^{S}_{\theta}$ is the encoder that we want to improve, we freeze teacher network $f^{T}_{\theta}$ which is a pre-trained encoder by training using an off-the-shelf method.
 Given an image x, the teacher network extracts its representation $h^{T}$, in which $h^{T}\in\mathbb{R}^{D_{T}}$ is the final feature without any classifier.
@@
 We use SimCLR and MoCo-v2 as our contrastive learning methods, but any contrastive-based learning method can be adopted.
-3.2. Relational Self-supervised Distillation
+3.2 Relational Self-supervised Distillation
 Relational Self-supervised Distillation (RSD) is a method that transfers instance relations from a teacher network to a student network.
@@
 The key objective is to align the similarity score distribution generated by the student network with that of the teacher network.
-This alignment is achieved by minimizing the KL divergence between the similarity score distributions of the two networks, ensuring a close match in their respective interpretations of the instance similarities.
+This alignment is achieved by minimizing the KL-divergence between the similarity score distributions of the two networks, ensuring a close match in their respective interpretations of the instance similarities.
 We apply the same augmentations into an image $x_{i}$ as a batch and map the embeddings $h^{T}=f^{T}_{\theta}(\tilde{x_{i}})$ and $h^{S}=f^{S}_{\theta}(\tilde{x_{i}})$ where $h^{T},h^{S}\in\mathbb{R}^{D}$ and $f^{T}_{\theta}$ and $f^{S}_{\theta}$ denote the teacher and student network, respectively.
 We compute the cosine similarity between l2-normalized descriptors from the teacher network and the queue. For the teacher and student, the similarities are:
-(1) $sim(h^{T}_{i},q^{T}_{j})=\begin{bmatrix}h^{T}_{i}/\lVert h^{T}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$
-(2) $sim(h^{S}_{i},q^{T}_{j})=\begin{bmatrix}h^{S}_{i}/\lVert h^{S}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$
+$sim(h^{T}_{i},q^{T}_{j})=\begin{bmatrix}h^{T}_{i}/\lVert h^{T}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$ (1)
+$sim(h^{S}_{i},q^{T}_{j})=\begin{bmatrix}h^{S}_{i}/\lVert h^{S}_{i}\rVert_{2}\end{bmatrix}\cdot[q_{j}^{T}]^{\top}$ (2)
 where $sim$ represents the cosine similarity, and $q_{j}^{\top}$ denotes the transposed j-th component in the teacher’s queue $Q^{T}$ to compute cosine similarity.
 We calculate the probability of the i-th instance with the j-th component in the teacher’s queue.
-(3) $p^{T}_{i,j}=\frac{exp(sim(h^{T}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}exp(sim(h^{T}_{i},q^{T})/\tau^{T})},\quad where\;j\in[1\dots K]$
-(4) $p^{S}_{i,j}=\frac{exp(sim(h^{S}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q}exp(sim(h^{S}_{i},q^{T})/\tau^{S})},\quad where\;j\in[1\dots K]$
+$p^{T}_{i,j}=\frac{exp(sim(h^{T}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}exp(sim(h^{T}_{i},q^{T})/\tau^{T})},\;j\in[1\dots K]$ (3)
+$p^{S}_{i,j}=\frac{exp(sim(h^{S}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q}exp(sim(h^{S}_{i},q^{T})/\tau^{S})},\;j\in[1\dots K]$ (4)
 where $\tau^{T}$ and $\tau^{S}$ are the temperature parameters of the teacher and student networks, respectively.
@@
 $=\arg\min_{\theta_{S}}\sum_{i}^{N}\sum_{j}^{K}-\frac{\exp(\text{sim}(h^{t}_{i},q^{T}_{j})/\tau^{T})}{\sum_{q^{T}\sim Q}\exp(\text{sim}(h^{t}_{i},q^{T})/\tau^{T})}\cdot$
-(5) $\qquad\log\left(\frac{\exp(\text{sim}(h^{s}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q^{T}}\exp(\text{sim}(h^{s}_{i},q^{T})/\tau^{S})}\right)$
-3.3. Contrastive Learning for ICD
+$\qquad\log\left(\frac{\exp(\text{sim}(h^{s}_{i},q^{T}_{j})/\tau^{S})}{\sum_{q^{T}\sim Q^{T}}\exp(\text{sim}(h^{s}_{i},q^{T})/\tau^{S})}\right)$ (5)
+3.3 Contrastive Learning for ICD
 The size of the embeddings within the architecture of lightweight networks is generally small, and it is crucial to maintain this size during training while benefiting from knowledge distillation.
@@
 We use these final embeddings as our descriptors, which allows for simultaneous training using our relational distillation approach.
-We follow the MoCo-v2 (Chen et al., 2020a) training procedure by utilizing a negative instance queue generated by the momentum encoder.
-We employ the InfoNCE (Oord et al., 2018b) loss to guide contrastive learning, which is calculated as follows:
-(6) $\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z^{S}_{i},z^{S^{\prime}}_{i})/\tau)}{\sum_{q^{S}\sim Q^{S}}\exp(\text{sim}(z^{S}_{i},q^{S})/\tau)}$
+We follow the MoCo-v2 [7] training procedure by utilizing a negative instance queue generated by the momentum encoder.
+We employ the InfoNCE [26] loss to guide contrastive learning as follows:
+$\mathcal{L}_{con}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\exp(\text{sim}(z^{S}_{i},z^{S^{\prime}}_{i})/\tau)}{\sum_{q^{S}\sim Q^{S}}\exp(\text{sim}(z^{S}_{i},q^{S})/\tau)}$ (6)
 For each pair index $i$, $z^{S}_{i}$, and $z^{S^{\prime}}_{i}$ are the representations of the two augmented views of the same image (i.e., a positive pair).
@@
 The denominator sums the exponentiated similarities of $z^{S}_{i}$ with all negative descriptors $q^{S}$ in all students’ queue $Q^{S}$.
-3.4. The Hard Negative Loss
+3.4 The Hard Negative Loss
 In the copy detection scenario, the primary challenges for training arise from the hard negative pairs and the fact that the use of small descriptors renders the environment susceptible to dimensional collapse.
-To address this, we employ the Hard Negative (HN) loss (Kordopatis-Zilos et al., 2023) to finely control the entropy and minimize the hard negative pairs.
+To address this, we employ the Hard Negative (HN) loss [21] to finely control the entropy and minimize the hard negative pairs.
 We also introduce an additional term that ensures that a hard negative pair can push their descriptors apart.
-(7) $\mathcal{L}_{hn}=-\frac{1}{N}\log\sum_{i=1}^{N}max_{j\in n_{i,j}}(1-S_{i,j})$
+$\mathcal{L}_{hn}=-\frac{1}{N}\log\sum_{i=1}^{N}max_{j\in n_{i,j}}(1-S_{i,j})$ (7)
 $n_{i,j}$ denotes the row indices of the negative pairs for the i-th row of the similarity matrix.
@@
 The final objective can be expressed as:
-(8) $\mathcal{L}_{RDCD}=\lambda_{rel}\mathcal{L}_{rel}+\lambda_{con}\mathcal{L}_{con}+\lambda_{hn}\mathcal{L}_{hn}$
-4. Experiments
-4.1. Dataset
-DISC2021 is a dataset consisting of training, reference and query images, that is used for the Image Similarity Challenge (Douze et al., 2021b).
+$\mathcal{L}_{RDCD}=\lambda_{rel}\mathcal{L}_{rel}+\lambda_{con}\mathcal{L}_{con}+\lambda_{hn}\mathcal{L}_{hn}$ (8)
+4 Experiments
+4.1 Dataset
+DISC2021 is a dataset consisting of training, reference and query images, that is used for the Image Similarity Challenge [8].
 The query image set contains strong auto-augmented and human-augmented images, and distractive images that may not appear in the reference set.
 The training set is designed to train a model without any labels, and the reference set is used for searching the background as a database.
-Copydays (Jegou et al., 2008) is a dataset designed to detect copied content, that consists of 157 original images and 2,000 queries with cropping, jpeg compressing and strong augmentations.
-We add 10k distractors from YFCC100M (Thomee et al., 2016), a common practice (Caron et al., 2021; Berman et al., 2019) that is known as CD10K.
+Copydays [17] is a dataset designed to detect copied content, that consists of 157 original images and 2,000 queries with cropping, jpeg compressing and strong augmentations.
+We add 10k distractors from YFCC100M [40], a common practice [5, 3] known as CD10K.
 We evaluate our method based on the mean average precision(mAP) and micro average precision($\mu AP$) for the strongly transformed copies.
-NDEC (Wang et al., 2023) dataset incorporates HN distractors, making it more advanced than conventional ICD datasets.
+NDEC [44] dataset incorporates HN distractors, making it more advanced than conventional ICD datasets.
 The basic data is derived from DISC2021 while the HN images are sourced from OpenImage.
@@
 The test data contains 49,252 query images and 1,000,000 reference images, with 24,252 of the query images being HN.
-4.2. Evaluation Metrics
-To enable a fair comparison with state-of-the-art methods (Pizzi et al., 2022; Caron et al., 2021), we use $\mu AP$ (Perronnin et al., 2009), accuracy at 1, and recall at precision 90 as the evaluation metrics.
-The equation for $\mu AP$ is as follows:
-(9) $\mu AP=\sum\limits_{i=1}^{N}P(i)\triangle r(i)\in[0,1]$
+4.2 Evaluation Metrics
+To enable a fair comparison with state-of-the-art methods [31, 5], we use $\mu AP$ [30] as the evaluation metric. The equation for $\mu AP$ is as follows:
+$\mu AP=\sum\limits_{i=1}^{N}P(i)\triangle r(i)\in[0,1]$ (9)
 where $P(i)$ is the precision at position $i$ of the sorted list of pairs, $\triangle r(i)$ is the difference in the recall between position i and i-1, and N is the total number of returned pairs for all queries.
 Any detected pair for a distractor query will decrease the average precision, thus all queries are evaluated together by merging the returned pairs for all queries, sorting them by confidence, and generating a single precision-recall curve.
-This is different from mAP, also known as macro-AP (Perronnin et al., 2009), where the average precision is computed separately per query and then averaged over all queries.
+This is different from mAP, also known as macro-AP [30], where the average precision is computed separately per query and then averaged over all queries.
 $\mu AP$ considers confidence values to capture all queries, while mAP only considers the query in the ground truth.
@@
 The former is more appropriate for ICD, while the latter is typically used in retrieval tasks.
-[…]
-Method Network Size $\mu$AP $\mu$APSN
-Multigrain RN-50 1500 16.5 36.5
-Multigrain RN-50 2048 20.5 41.7
-DINO ViT-B/8 1536 32.6 54.4
-DINO ViT-B/16 1536 32.2 53.8
-SSCD RN-50 512 61.5 72.5
-SSCD ResNeXt-101 1024 63.7 75.3
+We additionally evaluate with $\mu AP_{SN}$.
+Score normalization is a post-processing technique commonly used in image search to achieve a more uniform distribution.
+It adjusts the scores based on an estimate of the density around the query and normalizes them using background images that were not involved in training.
+Table 1: Results for the DISC2021 showing different methods, network architectures (RN-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (micro Average Precision $\mu$AP and micro Average Precision with score normalization $\mu$APSN).
+* means our implementation.
+$+$ used as teacher model in our experiments.
+Method Network $dim$ $\mu$AP $\mu$APSN
+Multigrain [3] RN-50 1500 16.5 36.5
+Multigrain [3] RN-50 2048 20.5 41.7
+DINO+ [5] ViT-B/8 1536 32.6 54.4
+DINO [5] ViT-B/16 1536 32.2 53.8
+SSCD+ [31] RN-50 512 61.5 72.5
+SSCD+ [31] ResNeXt-101 1024 63.7 75.3
 using lightweight
@@
 RDCD EN-B0 256 52.7 65.7
-4.3. Training Implementation
+4.3 Training Implementation
 We train all networks using the pretraining parameters from ImageNet-1K.
 Batch normalization statistics are synchronized across all GPUs.
-We adhere to the hyperparameters outlined in SimCLR (Chen et al., 2020b), MoCo-v2 (Chen et al., 2020a) and SSCD (Pizzi et al., 2022).
-The batch size is set at N = 256, with a resolution of 224 x 224 for training and 288 x 288 for inference.
+We adhere to the hyperparameters outlined in SimCLR [6], MoCo-v2 [7] and SSCD [31].
+The batch size is set at N = 256, with a resolution of 224 x 224 for training and 288 x 288 for evaluation.
 We use a learning rate of 1.0 and a weight decay of $10^{-6}$.
@@
 For our knowledge distillation process, we set the temperature at 0.04 for the teacher and 0.07 for the student.
-All experiments are conducted on an NVIDIA A100 80GB GPU.
+All experiments are conducted on 1 NVIDIA A100 80GB GPU.
 To ensure reproducibility, we maintain consistent seed values throughout all of the experiments.
-4.4. Results
+We further discuss the practicality on RDCD in Appendix E.
+4.4 Results
 Table 1 presents the results for the DISC2021 dataset.
-The first section of the table summarizes the performance of different methods with varying descriptor sizes, while the second section describes the performance of our baseline SSCD (Pizzi et al., 2022) strategy using a lightweight network with descriptor sizes of 64, 128, and 256.
-When training using a lightweight network, we do not use a mix-up for the fair comparison of our methods.
+The first section of the table summarizes the performance of different methods with varying descriptor sizes, while the second section presents the performance of our baseline SSCD [31] strategy using a lightweight network with descriptor sizes of 64, 128, and 256.
+When training using a lightweight network, we do not use mix-up to ensure a fair comparison of our methods.
 The third section shows the results of the proposed RDCD approach.
 When compared with the methods outlined in the first section, our RDCD method achieves competitive performance despite using significantly smaller descriptors and a more compact network size.
-It is noteworthy that the RDCD method, with a descriptor size of 64, achieves a $\mu$APSN of 53.5, which is comparable to the DINO method that utilizes a ViT-B/16 network with a descriptor size of 1536, yielding a $\mu$APSN of 53.8.
+It is noteworthy that the RDCD method, with a descriptor size of 64, achieves a $\mu$APSN of 53.5, which is comparable to the DINO method, which utilizes a ViT-B/16 network with a descriptor size of 1536, yielding a $\mu$APSN of 53.8.
 This comparison is significant, as the descriptor size of RDCD is 24 times smaller.
-When evaluated against an equivalent lightweight architecture, RDCD consistently outperforms SSCD across all three dimensions assessed.
-We employ the SSCD model RN-50, which has a $\mu$APSN of 72.5, as the teacher network to provide knowledge to the lightweight network.
+When evaluated with an equivalent lightweight architecture, RDCD consistently outperforms SSCD across all three dimensions assessed.
+We use the SSCD model RN-50, which has a $\mu$APSN of 72.5, as the teacher network to transfer knowledge to the lightweight network.
 When we apply our methods to EfficientNet-B0, RDCD outperforms SSCD.
@@
 This indicates that our method is capable of delivering strong results even with smaller descriptors.
-RDCD also outperforms SSCD in terms of mAP using the CD10K dataset (Table 2).
+RDCD also outperforms SSCD using the CD10K dataset in Table 2.
 Specifically, when using EfficientNet-B0 with a descriptor size of 128, RDCD achieves an mAP of 79.2, which is significantly higher than the best result for SSCD.
-With a descriptor size of 256, the performance improvement is even more pronounced, with RDCD reaching a mAP of 81.4.
+With a descriptor size of 256, the performance improvement becomes more significant, achieving an mAP of 81.4.
 For the NDEC dataset, Table 3 shows that our method shows competitive performance compared to the teacher network SSCD-RN50, which achieves a $\mu$APSN of 23.3 with a descriptor size of 512.
-Our proposed RDCD approach, employing EN-B0 network with descriptor sizes of 64, 128, and 256 achieves a $\mu$APSN of 17.3, 19.5, and 21.3, respectively.
+Our proposed RDCD approach, employing EN-B0 network with a descriptor size of 64, 128, and 256 achieves a $\mu$APSN of 17.3, 19.5, and 21.3, respectively.
 This demonstrates that across all compact descriptor sizes, RDCD surpasses the performance of the previous SSCD when using the same lightweight architecture.
-Table 2.
-Results on the CD10K (Copydays+ 10k distractors).
-We compare different methods, network architectures (ResNet-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (mean Average Precision (mAP) and micro Average Precision ($\mu$AP).
+Table 2: Results on the CD10K (Copydays+ 10k distractors).
+We compare different methods, network architectures (RN-50, ViT-B/8, ViT-B/16, ResNeXt-101, EfficientNet-B0), model sizes, and performance metrics (mean Average Precision (mAP) and micro Average Precision ($\mu$AP).
 * means our implementation.
-Method Network Size mAP $\mu$AP
-Multigrain RN-50 1500 82.3 77.3
-DINO ViT-B/8 1536 85.3 91.7
-DINO ViT-B/16 1536 80.7 88.7
-SSCD RN-50 512 85.0 97.9
-SSCD ResNext-101 1024 91.9 96.5
-using lightweight
+Method Network $dim$ mAP $\mu$AP
+Multigrain [3] RN-50 1500 82.3 77.3
+DINO [5] ViT-B/8 1536 85.3 91.7
+DINO [5] ViT-B/16 1536 80.7 88.7
+SSCD [31] RN-50 512 85.0 97.9
+SSCD [31] ResNeXt-101 1024 91.9 96.5
 SSCD EN-B0 64 64.1 95.6
@@
 RDCD EN-B0 256 81.4 95.4
-Table 3. Comparison of NDEC results. * means our implementation.
-Method Network Size $\mu$AP $\mu$APSN
-DINO ViT-B/8 1536 0.08 11.4
-DINO ViT-B/16 1536 0.09 13.1
-SSCD RN-50 512 21.2 23.3
+Table 3: Comparison of NDEC results. * means our implementation.
+Method Network $dim$ $\mu$AP $\mu$APSN
+DINO [5] ViT-B/8 1536 16.2 22.8
+DINO [5] ViT-B/16 1536 18.1 26.2
+SSCD [31] RN-50 512 42.4 46.6
 using lightweight
-SSCD EN-B0 64 14.4 16.4
-SSCD EN-B0 128 16.5 18.7
-SSCD EN-B0 256 17.5 20.1
-SSCD EN-B0 512 17.3 20.4
+SSCD EN-B0 64 28.9 32.8
+SSCD EN-B0 128 33.0 37.5
+SSCD EN-B0 256 35.0 40.1
+SSCD EN-B0 512 34.5 40.8
 ours
-RDCD EN-B0 64 15.7 17.3
-RDCD EN-B0 128 17.5 19.5
-RDCD EN-B0 256 18.8 21.3
-5. Discussion
-5.1. Effect of RSD and HN Loss
+RDCD EN-B0 64 31.4 34.7
+RDCD EN-B0 128 34.9 38.9
+RDCD EN-B0 256 37.5 42.5
+5 Discussion
+5.1 Effect of RSD with HN Loss
 Our method employs an additional projector to reduce the descriptor size, which creates a network with a two-layer MLP.
 This configuration can lead to implicit regularization, due to the interaction between the weight matrices across different layers.
-Implicit regularization typically restricts the ability of a network to learn diverse features, leading to dimensional collapse with contrastive SSL (Jing et al., 2021).
-However, when applying HN loss together with contrastive learning, we show that dimensional collapse does not occur even with the use of an MLP, thus reducing the impact of implicit regularization.
-Figure 3. Log of singular values for a descriptor size of 256, with and without HN loss.
+Implicit regularization typically restricts the ability of a network to learn diverse features, leading to dimensional collapse with contrastive SSL [18].
+However, when applying HN loss together with RSD, we show that dimensional collapse does not occur even with the use of an MLP, thus reducing the impact of implicit regularization.
+Figure 3: Log of singular values for a descriptor size of 256, with and without HN loss.
 To verify this, we calculate the singular values of the final descriptor of a model with a descriptor size of 256 under four conditions: with the use of RSD, FKD and with and without the use of HN loss.
-We compute the descriptors for 50k queries from the DISC21 dataset, calculate their covariance, and perform singular value decomposition on the covariance matrix.
+We compute the descriptors for 50k queries from the DISC2021 dataset, calculate their covariance, and perform singular value decomposition on the covariance matrix.
 Subsequently, we extract the singular values and apply a logarithmic operation to them.
@@
 It can also be observed that applying HN loss with FKD does not have a significant effect.
-During the training, we visualize the similarity of the hard negative for both the MoCo-v2 baseline and HN loss, implemented with either RSD or FKD, using descriptor sizes of 128 or 256, as shown in Figure 4.
-FKD transfers knowledge by aligning embeddings of the student model with those of the teacher model using mean squared error (MSE) loss.
-As Figure 4 shows, using FKD with HN loss fails to reduce the similarity of the nearest negatives, whereas using RSD with HN loss gradually reduces the nearest negative similarity.
-We further visualize the difference between positive similarity and hard negative similarity across multiple dimensions experimented in MoCo-v2 — RSD environment, comparing scenarios with and without the application of HN loss.
-We analyze a sample of 5,000 distinct queries extracted from the DISC21 dataset, which are separate from the queries used in the standard evaluation process for generality, to calculate and assess the differences in the similarity between positive and hard negative samples.
+We further visualize the difference between positive similarity and hard negative similarity across multiple dimensions experimented in MoCo-v2 |||| RSD environment, comparing scenarios with and without the application of HN loss.
+We analyze a sample of 5,000 distinct queries extracted from the DISC2021 dataset, which are separate from the queries used in the standard evaluation process for generality, to calculate and assess the differences in the similarity between positive and hard negative samples.
 As Figure 4 shows, across all descriptor sizes, the application of HN loss resulted in a more pronounced disparity compared to when it was omitted.
@@
 We conjecture that this trend may contribute to improved performance in higher-dimensional descriptors.
-Figure 4. Comparison of hard negative similarity of RSD and FKD.
 (a) 64 descriptors
@@
 (c) 256 descriptors
-Figure 5. Comparison of the difference in similarity between positives and nearest negatives with and without the use of HN Loss.
-Table 4. Comparison of other distillation. All methods are trained with EfficientNet-B0 network. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
-128 descriptors
-Hard Negative Loss No Yes
-Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
-SSCD 43.5 56.2 - -
-SimCLR 46.2 58.7 48.4 59.9
-47.4 59.6 49.6 61.4
-47.5 59.6 50.8 61.5
-MoCo-v2 40.3 57.5 44.8 59.1
-47.1 60.7 50.0 61.1
-46.2 59.3 49.6 61.0
-256 descriptors
-Hard Negative Loss No Yes
-Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
-SSCD 46.0 59.8 - -
-SimCLR 51.7 64.0 51.4 64.5
-52.2 65.6 52.7 65.8
-52.4 65.5 49.9 64.6
-MoCo-v2 41.3 56.6 46.1 60.1
-47.2 61.0 52.9 65.7
-47.3 60.8 53.2 65.9
-5.2. Comparison with Other Forms of Distillation
-To further demonstrate the efficacy of RSD, we conduct a comparative analysis with FKD by 1) employing FKD alone, 2) employing RSD alone, and 3) integrating both FKD and RSD (Table 13).
-We also employ SimCLR-style contrastive learning.
-When FKD and RSD are combined, the best performance is observed in several cases, suggesting that combining both distillation methods can further improve performance.
-Nevertheless, RSD consistently outperforms FKD in all cases, with RSD alone achieving performance on par with the combined use of both distillation methods.
-This indicates that RSD is sufficiently robust, and FKD does not significantly alter the outcome.
-It confirms that RSD can operate alone to enhance the performance of compact descriptors.
-Additional experiments with other distillation methods are presented in Appendix A.
-5.3. Effect of Different Architectures
-We demonstrate the effectiveness of our proposed approach in scenarios where the teacher networks belong to different architectural families (ViT or CNN), with student models also selected from these two distinct architectures.
-For the ViT, we use DINO as the teacher, and for the CNNs, we use SSCD-RN50 and SSCD-ResNeXt-101.
-All models are trained on SSCD-RN50 hyperparameter settings.
-With DINO as the teacher model, we follow the ICD settings presented in (Caron et al., 2021) for the DINO baseline.
-Consequently, we distill a descriptor size of 1536 from the teacher (concatenation of 768 class tokens and 768 GeM pooled patch tokens).
-In the case of ResNeXt-101 teacher model, we distill the final target descriptor with a size of 1024.
-Table 5 shows that our method is highly effective for the different architectures, highlighting its adaptability and robustness.
-Table 5. DISC evaluation with different architectures. All experiments are performed with a descriptor size of 128.
-T S Eff-B0 FastViT-T12 Mob-V3
-$\mu$AP $\mu$APSN $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
-DINO-ViT-B/8 32.1 52.1 28.1 52.2 39.6 54.7
-SSCD-RN-50 50.0 61.1 48.5 61.1 47.3 58.8
-SSCD-ResNeXt-101 44.9 59.9 42.8 57.4 41.1 56.6
-5.4. Ablation Study
+Figure 4: Comparison of the difference in similarity between positives and nearest negatives with and without the use of HN Loss.
+5.2 Effect of Hard Negative Loss on Uniformity
+A uniform distribution of descriptors represents semantically similar images far apart while placing copies of the same image close together.
+Both DINO [5] and SSCD [31] are used as teacher models; however, only SSCD uses entropy regularizer to achieve a uniform distribution.
+In general, features used during evaluation are called descriptors.
+We demonstrate that applying HN loss to the final feature used as a descriptor results in a more uniform distribution.
+This effect occurs regardless of whether the teacher was trained with or without HN loss.
+Figure 5 visually supports this observation.
+b) shows the distribution of a student model distilled from a non-uniform teacher and trained with HN loss, while c) displays the distribution of a student model distilled from a uniform teacher, also trained with HN loss.
+Both distributions are notably more uniform compared to the distribution of the DINO teacher in a).
+We further analyze the intermediate feature, which is directly distilled by the teacher model, and show that when using RSD, the uniformity of the teacher’s embedding is transferred to the student’s embedding.
+d) shows the result of distillation from a non-uniform teacher DINO, while e) shows the result of distillation from a uniform teacher SSCD.
+Consequently, if the teacher’s embedding has a uniform distribution in the embedding space, applying HN loss to the final feature may yield limited additional uniformity.
+Additional experiments with intermediate features are illustrated in Appendix B.
+Figure 5: T-SNE visualization of descriptors from different model configurations: a) DINO baseline, b) DINO-MoCo-v2 w/ HN loss, c) SSCD-MoCo-v2 w/ HN loss.
+d) DINO-MoCo-v2 w/o HN loss, e) SSCD-MoCo-v2 w/o HN loss.
+In a),b) and c), the final features are used as descriptors whereas in d) and e), intermediate features are used.
+We randomly select 9 classes and visualize 50 images from each class from ImageNet.
+5.3 Ablation Study
 Effectiveness of Each Loss Term.
@@
 A low RPR indicates dimensional collapse, and RPR value of 1 indicates that the descriptor makes full use of all dimensions.
-All models are trained on the DISC21 dataset with 100 epochs and SSCD-RN50-w/mixup is employed as the teacher model.
+All models are trained on the DISC2021 dataset with 100 epochs and SSCD-RN50-w/mixup is employed as the teacher model.
 Our result shows that with the addition of RSD loss and HN loss, the performance improves substantially, which indicates that RSD and HN loss can benefit the performance of RDCD.
-Table 6. Ablation study on the loss terms in RDCD. Descriptor sizes of 64, 128, and 256 were examined, and a consistent MLP configuration was maintained across all experiments for a fair comparison.
+Table 4: Ablation study on the loss terms in RDCD. Descriptor sizes of 64, 128, and 256 were examined, and a consistent MLP configuration was maintained across all experiments for a fair comparison.
 Loss MLP-d $\mu$AP $\mu$APSN
@@
 $\lambda_{rel}$ and $\lambda_{hn}$ are the weights of RSD loss and HN loss, respectively.
-First, we analyze the influence of contrastive loss and RSD loss (Table 7).
+First, we analyze the influence of contrastive loss and RSD loss, as shown in Table 5.
 As shown in the third row, a contrastive loss of 1 and RSD loss of 10 achieves the best performance, which indicates that knowledge distillation is essential for the model with better performance.
@@
 In the first section, where only contrastive loss and RSD loss are employed, we observe that RPR gradually increases incrementally with an increase in $\lambda_{rel}$.
-This suggests that the student model partially alleviates the dimensional collapse which is guided by a teacher model that has been trained with an entropy regularizer (SSCD).
+This suggests that the student model partially alleviates the dimensional collapse which is guided by a teacher model that has been trained with an entropy regularizer.
 Subsequently, in the second section, where HN loss is incorporated, the student model finally utilizes the full dimensions of the descriptors.
-Table 7. Ablation study to evaluate the impact of the loss ratio. All experiments are performed with a descriptor size of 256.
+Table 5: Ablation study to evaluate the impact of the loss ratio. All experiments are performed with a descriptor size of 256.
 $\lambda_{con}$ $\lambda_{rel}$ $\lambda_{hn}$ RPR $\mu$AP $\mu$APSN
@@
 1 10 above 5 - collapse collapse
-We next conduct a step-by-step experiment with DINO as the teacher model (Table 8).
+We next conduct a step-by-step experiment with DINO as the teacher model, as shown in Table 6.
 Our approach reduces the descriptor size by a factor of 12 but produces a performance comparable to that of the teacher model.
-As illustrated in Discussion 5.1, dimensional collapse occurs when HN loss is not employed.
+As illustrated in 5.1, dimensional collapse occurs when HN loss is not employed.
 Applying HN loss thus successfully improves performance and prevents dimensional collapse.
-Table 8. Ablation study of using DINO as teacher model.
-Method Network Size RPR $\mu$AP $\mu$APSN
-DINO (Teacher) ViT-B/8 1536 0.99 32.6 54.4
+Table 6: Ablation study of using DINO as teacher model.
+Method Network $dim$ RPR $\mu$AP $\mu$APSN
+DINO [5] (Teacher) ViT-B/8 1536 0.99 32.6 54.4
 RSD EN-B0 1536 0.88 10.0 35.2
-RSD — MoCo-v2 EN-B0 128 0.57 20.7 40.0
-RSD — MoCo-v2 — HN (RDCD) EN-B0 128 1.0 32.1 52.1
-6. Conclusion
+RSD |||| MoCo-v2 EN-B0 128 0.57 20.7 40.0
+RSD |||| MoCo-v2 |||| HN (RDCD) EN-B0 128 1.0 32.1 52.1
+6 Conclusion
 In this paper, we present RDCD, a novel method for training lightweight networks with compact descriptors for image copy detection in a self-supervised manner.
@@
 We believe that this approach offers significant advantages in search speed and scalability for multimedia applications.
+Acknowledgements
+This work was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2023-00224740, Development of technology to prevent and track the distribution of illegally filmed content).
 Appendix
-Appendix A Different Architectures
-To show the generality of our method, we adopt one lightweight model each from the most commonly used architectures in computer vision, CNN and ViT, conducting experiments with MobileNet-V3 and FastViT-T12.
+Appendix A Effect of Different Architectures
+A.1 Different Student
+To show the generality of RDCD, we adopt one lightweight model each from the most commonly used architectures in computer vision, CNN and ViT, conducting experiments with MobileNet-V3 and FastViT-T12.
 Our results are shown in Table 7, Table 8, Table 9.
-Table 9. Performance of DISC2021 with different student architectures. * means our implementation.
-Method Network Size $\mu$AP $\mu$APSN
+Specifically, using FastViT under the same conditions resulted in a $\mu$AP of 67.4, demonstrating the research potential for ViT student.
+Table 7: Performance of DISC2021 with different student architectures. * means our implementation.
+Method Network $dim$ $\mu$AP $\mu$APSN
 using lightweight
@@
 RDCD FastViT-T12 256 56.4 67.4
-Table 10. Performance of CD10K(Copydays+ 10k distractors) with different student architectures. * means our implementation.
-Method Network Size mAP $\mu$AP
+Table 8: Performance of CD10K(Copydays+ 10k distractors) with different student architectures. * means our implementation.
+Method Network $dim$ mAP $\mu$AP
 Multigrain RN-50 1500 82.3 77.3
@@
 RDCD FastViT-T12 256 79.3 97.0
-Table 11. Performance of NDEC with different student architectures. * means our implementation.
-Method Network Size $\mu$AP $\mu$APSN
+Table 9: Performance of NDEC with different student architectures. * means our implementation.
+Method Network $dim$ $\mu$AP $\mu$APSN
 DINO ViT-B/8 1536 16.2 22.8
@@
 RDCD FastViT-T12 256 39.1 43.2
-Appendix B Comparison with Barlow-Distill
-We introduce the RDCD method using the MoCo framework’s negative sample queue in our distillation process.
-Alternatively, there are self-supervised learning (SSL) methods that do not employ negative samples and instead focus solely on positive samples.
-Inspired by the Barlow Twins approach, we have integrated the capability of distillation in our process by generating positive pairs within a batch through augmentation and optimizing the cross-correlation matrix between these pairs.
-This method, we call Barlow-Distill, ensures that each feature operates as independently as possible from the others, thereby minimizing feature redundancy.
-This minimization ensures that each feature retains unique information, leading to a richer data representation.
-The Barlow-Distill approach distinctly avoids the use of negative samples, with all learning occurring through the adjustment of the correlation between positive pairs.
-This method is particularly effective when generating different transformations of the same image through augmentation techniques.
-Consequently, the model enhances the representation of the same object or scene in various forms.
-Barlow-Distill generalizes well by minimizing local features in two distinct augmented images.
-This process effectively captures higher levels of abstract information, ensuring that the latent representation retains sufficient information to learn useful representations for various downstream tasks.
-However, this concentration on high-level features may result in the loss of crucial low-level details, which can degrade performance.
-This loss is particularly problematic in image copy detection tasks, where missing fine-grained visual details can lead to degraded performance.
-In contrast, SSL methods utilizing contrastive learning could be more suitable for tasks such as image copy detection.
-This method, by clearly differentiating between positive and negative samples, enables the model to more effectively recognize and learn local detail and fine-grained visual differences.
-Contrastive learning enhances the model’s ability to distinguish between similar images, thus improving accuracy and efficiency in the image copy detection task.
-This enhancement is vital for capturing important local features in images, a crucial ability for image copy detection.
-According to our research results presented in Table 12, the RDCD model shows enhanced performance compared to Barlow-Distill.
-Table 12. Comparision with RDCD and Barlow-Distill
-Method Network Size $\mu$AP $\mu$APSN
-Barlow-Distill EN-B0 64 44.1 48.3
-Barlow-Distill EN-B0 128 51.9 59.0
-Barlow-Distill EN-B0 256 54.8 64.4
-RDCD EN-B0 64 43.9 53.5
-RDCD EN-B0 128 50.0 61.1
-RDCD EN-B0 256 52.7 65.7
-Appendix C Intermediate Descriptor
-In this study, we evaluate the performance of DISC2021 using an intermediate descriptor designed to align the descriptor size with that of the teacher model.
-We assess a 512-dimensional descriptor both with and without the application of HN loss, utilizing the EfficientNet-B0 network architecture.
-Our results (Table 13) show that the proposed RDCD method consistently outperforms SSCD across two projection sizes, 128 and 256.
-Furthermore, our analysis reveals that the performance gap between using and not using HN loss is negligible, and the intermediate descriptor fully exploits its 512-dimensional capacity without HN loss.
-This finding suggests that the student model is effectively trained under the guidance of a teacher model equipped with an entropy regularizer.
-Table 13. Comparison of descriptor size of 512. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
+A.2 Different Teacher
+We demonstrate the effectiveness of our proposed approach in scenarios where the teacher networks belong to different architectural families (ViT or CNN), with student models also selected from these two distinct architectures.
+For the ViT, we use DINO as the teacher, and for the CNNs, we use SSCD-RN50 and SSCD-ResNeXt-101.
+All models are trained on SSCD-RN50 hyperparameter settings.
+With DINO as the teacher model, we follow the ICD settings presented in [5] for the DINO baseline.
+Consequently, we distill a descriptor size of 1536 from the teacher (concatenation of 768 class tokens and 768 GeM pooled patch tokens).
+In the case of ResNeXt-101 teacher model, we distill the final target descriptor with a size of 1024.
+Table 10 shows that our method is highly effective for the different architectures, highlighting its adaptability and robustness.
+Table 10: DISC2021 evaluation with different architectures. All experiments are performed with a descriptor size of 128.
+T S EN-B0 FastViT-T12 Mob-V3
+$\mu$AP $\mu$APSN $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
+DINO-ViT-B/8 32.1 52.1 28.1 52.2 39.6 54.7
+SSCD-RN-50 50.0 61.1 48.5 61.1 47.3 58.8
+SSCD-ResNeXt-101 44.9 59.9 42.8 57.4 41.1 56.6
+Table 11: The following table demonstrates that when dimension reduction is applied to the descriptor through PCA whitening, our approach outperforms the existing baseline in terms of $\mu$AP performance.
+We do not apply Mixup.
+* means our implementation.
+Method Network $dim$ $\mu$AP $\mu$APSN $\mu$APSN256 $\mu$APSN128 $\mu$APSN64
+DINO ViT-B/8 1536 32.6 54.4 47.7 42.3 32.7
+DINO ViT-B/16 1536 32.2 53.8 47.4 42.0 32.0
+SSCD RN-50 512 43.6 72.5 66.0 56.3 38.3
+SSCD ResNext-101 1024 63.7 75.3 65.5 54.0 34.8
+using lightweight
+SSCD EN-B0 64 38.2 48.5 48.5 48.5 48.5
+SSCD EN-B0 128 43.5 56.2 56.2 56.2 44.4
+SSCD EN-B0 256 46.0 59.8 59.8 53.8 42.3
+SSCD EN-B0 512 43.6 61.1 58.0 52.0 41.9
+Ours
+RDCD EN-B0 64 43.9 53.5 53.5 53.5 53.5
+RDCD EN-B0 128 50.0 61.1 61.1 61.1 53.1
+RDCD EN-B0 256 52.7 65.7 65.7 61.5 52.2
+Table 12: Comparison of intermediate descriptor. We do not use HN loss for SSCD method as they use Koleo regularizer.
 128 projection
@@
 55.8 69.2 56.1 69.0
-Appendix D Qualitative Examples
+Appendix B Intermediate Descriptor
+In this study, we evaluate the performance of DISC2021 using an intermediate descriptor designed to align the descriptor size with that of the teacher model.
+We assess a 512-dimensional descriptor both with and without the application of HN loss, utilizing the EfficientNet-B0 network architecture.
+Our results (Table 12) show that the proposed RDCD method consistently outperforms SSCD across two projection sizes, 128 and 256.
+Furthermore, our analysis reveals that the performance gap between using and not using HN loss is negligible, and the intermediate descriptor fully exploits its 512-dimensional capacity without HN loss.
+This finding suggests that the student model is effectively trained under the guidance of a teacher model equipped with an entropy regularizer.
+Appendix C PCA
+The dimensionality of the descriptor plays a crucial role in optimizing the trade-off between matching time and accuracy during the matching phase of ICD.
+Previous approaches have employed PCA and PCA whitening to create compact descriptors.
+Following [31]The post-processing technique of PCA whitening is known for its efficacy in making the descriptor distribution more uniform.
+Our RDCD method, trained via deep learning, exhibits superior performance compared to the baseline, which utilizes large-sized descriptors that have undergone PCA whitening.
+These results are detailed in Table  11.
+Table 13: Comparison of other distillation. All methods are trained with EfficientNet-B0 network. We do not use Hard Negative loss for SSCD method as they use Koleo regularizer.
+128 descriptors
+Hard Negative Loss No Yes
+Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
+SSCD 43.5 56.2 - -
+SimCLR 46.2 58.7 48.4 59.9
+47.4 59.6 49.6 61.4
+47.5 59.6 50.8 61.5
+MoCo-v2 40.3 57.5 44.8 59.1
+47.1 60.7 50.0 61.1
+46.2 59.3 49.6 61.0
+256 descriptors
+Hard Negative Loss No Yes
+Method RSD FKD $\mu$AP $\mu$APSN $\mu$AP $\mu$APSN
+SSCD 46.0 59.8 - -
+SimCLR 51.7 64.0 51.4 64.5
+52.2 65.6 52.7 65.8
+52.4 65.5 49.9 64.6
+MoCo-v2 41.3 56.6 46.1 60.1
+47.2 61.0 52.9 65.7
+47.3 60.8 53.2 65.9
+Appendix D Comparison with Other Forms of Distillation
+To further demonstrate the efficacy of RSD, we conduct a comparative analysis with FKD by 1) employing FKD alone, 2) employing RSD alone, and 3) integrating both FKD and RSD in Table 13.
+We also employ SimCLR-style contrastive learning.
+When FKD and RSD are combined, the best performance is observed in several cases, suggesting that combining both distillation methods can further improve performance.
+Nevertheless, RSD consistently outperforms FKD in all cases, with RSD alone achieving performance on par with the combined use of both distillation methods.
+This indicates that RSD is sufficiently robust, and FKD does not significantly alter the outcome.
+It confirms that RSD can operate alone to enhance the performance of compact descriptors.
+Table 14: Ablation study on the computational cost during training.
+All experiments are conducted on a single A100 GPU using FP32 precision with EFF-B0 architecture and a descriptor size of 256.
+All metrics are calculated with a batch size of 256 except for FLOPs, which we measured with a batch size of 1.
+Method Contrastive Learning Distillation #Params FLOPs Time per Epoch VRAM
+(M) (G) (min) (G)
+SSCD SimCLR X 4.3 0.828 58.46 49.7
+- - O 29.3 8.732 40.06 28.1
+- SimCLR O 29.4 9.16 74.34 54.9
+RDCD MoCo-v2 O 34.2 9.9 51.25 29.8
+Table 15: Comparison of computational cost during evaluation, highlighting our method’s high efficiency. We evaluate images with 288x288 size. FLOPs is calculated with a batch size of 1.
+Method Architecture #Params FLOPs Throughput Search Time $dim$ $\mu$AP
+(M) (G) (image/sec)
+DINO ViT-B/8 85.204 133.699 87.76 37.5 1536 54.4
+SSCD RN-50 24.6 8.265 903.58 15.518 512 72.5
+RDCD EN-B0 4.8 0.829 1087.75 10.121 256 65.7
+Appendix E Practicality of RDCD
+We perform additional analysis of practicality on RDCD.
+To highlight the computational cost of each component, in Table 14 we conduct ablation experiments by removing contrastive learning and distillation.
+SSCD uses SimCLR without distillation.
+We found that our method requires 29.9 million more parameters than SSCD for training, primarily due to the large teacher network, which accounts for 24.0M parameters.
+Additionally, our method consumes a total of 9.9G of FLOPs: 8.3G for distillation and 1.6G for contrastive learning.
+Nevertheless, SSCD’s use of SimCLR leads to a proportional increase in VRAM usage with larger batch sizes.
+In our RDCD, we adopt MoCo-v2 which stores negative features in a queue, significantly reducing VRAM usage.
+By leveraging MoCo-v2, our method achieves 0.6 times lower VRAM usage compared to SSCD and also reduces training time.
+Furthermore, our final performance surpasses SSCD by 5.9.
+Our RDCD incurs higher computational cost during training but shows significant efficiency during evaluation.
+Specifically, RDCD requires only 4.8M parameters during evaluation, which includes an 0.5M additional parameters from the EFF-B0 architecture.
+In copy detection, the evaluation involves two stages: model inference and database search.
+High-speed search is crucial when retrieving from a large-scale image database in real-world applications.
+We evaluate search time using the DISC2021 dataset, which consists of 1M database images and 50k query images.
+As shown in Table 15, RDCD achieves superior efficiency across all metrics compared to previous methods.
+DINO and SSCD require 133.699G, 8.265G FLOPs, respectively, due to their large architectures, while RDCD operates with 161 times lower FLOPs than DINO and 9.9 times lower FLOPs than SSCD during model inference.
+Appendix F Qualitative Examples
 In Figure 6, we present the queries along with the top-2 results retrieved by both the RDCD and SSCD methods for cases where both methods accurately identify the ground truth.
@@
 For our matching evaluation, we employ a descriptor of size 256, utilizing the EN-B0 network architecture.
-Figure 6. Example of top-2 retrieval results from the DISC2021 dataset.
-Appendix E PCA
-The dimensionality of the descriptor plays a crucial role in optimizing the trade-off between matching time and accuracy during the matching phase of ICD.
-Previous approaches have employed PCA and PCA whitening to create compact descriptors.
-The post-processing technique of PCA whitening is known for its efficacy in making the descriptor distribution more uniform (Pizzi et al., 2022).
-Our RDCD method, trained via deep learning, exhibits superior performance compared to the baseline, which utilizes large-sized descriptors that have undergone PCA whitening.
-These results are detailed in Table  14.
-Table 14.
-The following table demonstrates that when dimension reduction is applied to the descriptor through PCA whitening, our approach outperforms the existing baseline in terms of $\mu$AP performance.
-We do not apply Mixup.
-* means our implementation.
-Method Network Descriptor Size $\mu$AP $\mu$APSN $\mu$APSN256 $\mu$APSN128 $\mu$APSN64
-DINO ViT-B/8 1536 32.6 54.4 47.7 42.3 32.7
-DINO ViT-B/16 1536 32.2 53.8 47.4 42.0 32.0
-SSCD RN-50 512 43.6 72.5 66.0 56.3 38.3
-SSCD ResNext-101 1024 63.7 75.3 65.5 54.0 34.8
-using lightweight
-SSCD EN-B0 64 38.2 48.5 48.5 48.5 48.5
-SSCD EN-B0 128 43.5 56.2 56.2 56.2 44.4
-SSCD EN-B0 256 46.0 59.8 59.8 53.8 42.3
-SSCD EN-B0 512 43.6 61.1 58.0 52.0 41.9
-Ours
-RDCD EN-B0 64 43.9 53.5 53.5 53.5 53.5
-RDCD EN-B0 128 50.0 61.1 61.1 61.1 53.1
-RDCD EN-B0 256 52.7 65.7 65.7 61.5 52.2
+Figure 6: Example of top-2 retrieval results from the DISC2021 dataset.